### 09 Jan 2024 v1

# Phase-shifted remote photoplethysmography for estimating heart rate and blood pressure from facial video

Abstract
Human health can be critically affected by cardiovascular diseases, such as hypertension, arrhythmias, and stroke. Heart rate and blood pressure are important biometric information for the monitoring of cardiovascular system and early diagnosis of cardiovascular diseases. Existing methods for estimating the heart rate are based on electrocardiography and photoplethyomography, which require contacting the sensor to the skin surface. Moreover, catheter and cuff-based methods for measuring blood pressure cause inconvenience and have limited applicability. Therefore, in this thesis, we propose a vision-based method for estimating the heart rate and blood pressure. This thesis proposes a 2-stage deep learning framework consisting of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). In the first stage, DRP-Net infers remote photoplethysmography (rPPG) signals for the acral and facial regions, and these phase-shifted rPPG signals are utilized to estimate the heart rate. In the second stage, BBP-Net integrates temporal features and analyzes phase discrepancy between the acral and facial rPPG signals to estimate SBP and DBP values. To improve the accuracy of estimating the heart rate, we employed a data augmentation method based on a frame interpolation model. Moreover, we designed BBP-Net to infer blood pressure within a predefined range by incorporating a scaled sigmoid function. To demonstrate the effectiveness of the proposed method, experiments were conducted on the MMSE-HR and V4V datasets. Our method resulted in estimating the heart rate with the mean absolute error (MAE) of 1.78 BPM, reducing the MAE by 34.31 % compared to the recent method, on the MMSE-HR dataset. The MAE for estimating the systolic blood pressure (SBP) and diastolic blood pressure (DBP) were 10.19 mmHg and 7.09 mmHg. On the V4V dataset, the MAE for the heart rate, SBP, and DBP were 3.83 BPM, 13.64 mmHg, and 9.4 mmHg, respectively. We expect that the proposed method will contribute to the advancement of sensing technology for the monitoring of human health.
keywords:
Computer vision, deep learning, physiological measurement, remote photoplethysmography, heart rate, blood pressure
journal: ArXiv preprint
1 Introduction
The cardiovascular system consists of the heart and blood vessels, and it circulates blood throughout the body, delivering oxygen and nutrients to tissues while removing waste substances. Abnormalities in the cardiovascular system can lead to various cardiovascular diseases such as cardiomyopathy, hypertensive heart disease, arrhythmias, with potentially severe implications for overall health. Moreover, cardiovascular diseases associated with blood vessels, such as hypertension and stroke, are contribute to increasing the global mortality rate (Gaidai et al., 2023). However, these diseases can be detected at an early stage through the monitoring of the heart rate and blood pressure using healthcare devices such as smartwatches. For example, arrhythmias caused by abnormalities in the sinoatrial node can be prevented through the measurement of bradycardia, tachyarrhythmia, and irregular heart rates. Additionally, consistent readings of systolic blood pressure (SBP) above 120 mmHg and diastolic blood pressure (DBP) above 90 mmHg may indicate the potential presence of hypertension. Recently, interest in cardiovascular diseases has led to growing attention on research for healthcare services (Etemadi et al., 2023; Aggarwal & Pandey, 2023) and estimation of biometric information (Alnaggar et al., 2023; Gupta et al., 2023).
Heart rate and blood pressure are main biometric information for the monitoring of cardiovascular diseases (Perret-Guillaume et al., 2009; Fuchs & Whelton, 2020), and these information can be measured by utilizing various physiological sensors and medical equipment. Heart rate is commonly measured by using electrocardiography (ECG) and photoplethysmography (PPG). ECG uses electrodes attached to the body to record the electrical activity caused by contractions and relaxations of the heart. PPG is a non-invasive method that uses a light source and photodetector at the skin surface to measure the volumetric variations of blood in micro vessels. The heart rate is measured by computing peak-to-peak intervals in the time domain or by analyzing the power spectral density in the frequency domain from the ECG or PPG signals. On the other hand, blood pressure can be measured based on an oscillometric method, which records the magnitude of pressure oscillations using a cuff, or an invasive catheter-based approach. The catheter-based method involves the direct insertion of a sensor into an artery to measure continuous arterial blood pressure (ABP). Peak and valley values of the ABP signal are utilized as SBP and DBP, respectively. The blood pressure can also be estimated from pulse transit time (PTT) measured by ECG or PPG sensors at the heart and acral sites. PTT is the time taking for a blood pulse wave to travel from the heart to an acral site, and it known that PTT is closely correlated with blood pressure (Geddes et al., 1981). Although there have been proposed BP estimation methods by analyzing PTT from ECG or PPG signals (Barvik et al., 2021; Ganti et al., 2020), these approaches have the intrinsic limitation of requiring skin contact.
Recently, camera sensors have been utilized to obtain physiological information, and it is called remote photoplethysmography (rPPG). From a facial video, the rPPG technique extracts subtle variations in skin color induced by cardiovascular pulse. The extracted temporal changes in pixel intensities are then transformed into continuous waveforms analogous to conventional PPG signals. The precision of the rPPG technique can be critically influenced by many factors such as ambient light conditions, motion artifacts, and different skin tones. Despite these challenges, rPPG receives much attention due to the non-contact and non-invasive attributes that facilitate remote monitoring of biometric information.
Deep learning methods have been utilized to extract rPPG signals and to estimate the heart rate (Chen & McDuff, 2018; Yu et al., 2022; Ouzar et al., 2023). Additionally, various datasets (Bobbia et al., 2019; Niu et al., 2019; Soleymani et al., 2011; Li et al., 2018; Stricker et al., 2014; Zhang et al., 2016b; Revanur et al., 2021) have been released for the development and evaluation of these deep learning algorithms. PPG signals in the existing datasets were collected at acral sites such as fingertips, and therefore, there exist temporal discrepancies between the pulse waves in the PPG signals and facial videos. However, most previous methods ignore the time shift between the two modalities and adopt a time domain loss between PPG signals measured at an acral site and rPPG signals extracted from facial videos. For example, Lu et al. (2021) and Yu et al. (2022) employed negative Pearson correlation loss to measure the similarity between the PPG signals and estimated rPPG signals in the time domain. In this paper, we introduced two types of phase-shifted rPPG signals, referred to as acral rPPG and facial rPPG, to consider the temporal discrepancy between facial videos and PPG signals measured at an acral site. The acral rPPG is guided by a time domain loss function, while the facial rPPG is guided by frequency domain features which have temporally global characteristics.
On the other hand, blood pressure can be estimated from PPG signals, both ECG and PPG signals, or facial videos. While PPG-based methods are widely utilized in wearable devices, this approach has the limitation of requiring physical contact to obtain PPG signals. Although the approach of utilizing both ECG and PPG signals can measure blood pressure based on PTT information, it also has the disadvantage of requiring contact with ECG and PPG sensors. Recently, camera-based methods have been proposed to estimate blood pressure in a non-contact manner by utilizing spatiotemporal features in facial videos. However, there is a performance gap between the camera-based and PPG-based methods. The objective of this study is to improve the performance of the camera-based approach by utilizing phase-shifted information in the acral and facial rPPG signals.
This paper proposes a 2-stage deep learning pipeline consisting of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net). In the first stage, DRP-Net infers the acral and facial rPPG signals, and it is used to estimate the heart rate. In the second stage, BBP-Net integrates temporal features based on a multi-scale fusion (MSF) module and analyzes phase discrepancy between the acral and facial rPPG signals to estimate SBP and DBP values. To improve the precision of blood pressure estimation, a scaled sigmoid function is employed in the BBP-Net to constrain the estimate into a certain range. Experiments were conducted on the MMSE-HR database (Zhang et al., 2016b) and V4V database (Revanur et al., 2021) to demonstrate the effectiveness of the proposed method for estimating heart rate and blood pressure from facial videos. The main contributions of this paper can be summarized as follows.
  • 1.
    We propose a novel 2-stage deep learning framework consisting of DRP-Net and BBP-Net networks.
  • 2.
    We employ a frame interpolation algorithm for the augmentation of bradycardia and pyknocardia video clips.
  • 3.
    We introduce the concept of phase-shifted rPPG for considering the temporal discrepancy between pulse waves in facial videos and PPG signals.
  • 4.
    We propose a novel loss function and employ a scaled sigmoid layer to improve the accuracy of heart rate and blood pressure.
  • 5.
    Our proposed method achieved the state-of-the-art performance for estimating heart rate and blood pressure on the MMSE-HR and V4V database.
The rest of this paper is organized as follows. Section 2 presents related work. Section 3.1 presents overall pipeline of the proposed method. Section 3.2 and Section 3.3 explain details about preprocessing and data augmentation process. Section 3.4, Section 3.5, and Section 3.6 present the DRP-Net, BBP-Net, and loss functions, respectively. Sections 4 and Section 5 present experimental results and conclusion.
2 Related work
2.1 Estimation of rPPG signals from facial videos
Image processing techniques and mathematical models have been utilized to analyze subtle variations in color intensities and to estimate rPPG signals. Poh et al. (2010) computed pulse waves by averaging on RGB color channels, and independent component analysis was conducted to extract rPPG signals. Lewandowska et al. (2011) estimated rPPG signals by conducting a channel selection process and principal component analysis on color intensities in a forehead region. De Haan & Jeanne (2013) proposed a chrominance-based method, called CHROM, to improve the performance of rPPG against motion artifacts. Wang et al. (2016) proposed the POS algorithm, which analyzes the plane orthogonal to the skin tone in the normalized RGB space. Meanwhile, with the success of deep learning methods for computer vision tasks, data driven approaches started to be applied in the estimation of rPPG signals.
The deep learning-based rPPG estimation technique focuses on learning trends in PPG that are robust to motion and external brightness changes from video sequence data with spatiotemporal characteristics. Chen & McDuff (2018) introduced a convolutional attention network (CAN) that employs a Siamese-structured 2 dimensional CNN model. This model takes both a difference map of adjacent viewpoints and the original viewpoint image. Spatial attentions are utilized to emphasize skin regions in the difference images, effectively integrating the models. Nowara et al. (2021) aimed to estimate spatial noise factors including motion and illumination using the inverse of spatial attentions of the CAN model. They integrated the LSTM to enhance spatial denoising and improve temporal robustness in PPG. Yu et al. (2019) proposed PhysNet, which employs a 3D CNN-based encoder to extract spatiotemporal features and a decoder along the time axis to estimate rPPG. They also introduced a loss function based on the Pearson correlation coefficient. Yu et al. (2022) proposed a loss function based on temporal difference convolution (TDC) using neighboring frame differences of a 4 dimensional feature map as query and key, and KL divergence of periodic component distributions. Yu et al. (2023) is a SlowFast Network for temporal improvement of the estimated signal, which has shown promising heart rate estimation accuracy even in cross-domain. However, existing studies primarily focuses on providing performance results of heart rate estimation using the power spectrum of rPPG, and less consideration is given to the blood pressure estimation steps based on rPPG.
2.2 Deep learning methods for estimating blood pressure
For non-invasive blood pressure estimation, researchers have attempted to develop deep learning algorithms utilizing physiological signal measured from various sensors. Miao et al. (2020) proposed a deep learning model based on ResNet and LSTM for continuous blood pressure estimation from single channel ECG. Panwar et al. (2020) introduced PP-Net, which consists of a 1D convolution and LSTM, to estimate SBP, DBP, and heart rate from PPG signals. Huang et al. (2022) utilized PTT information for blood pressure estimation and proposed a deep learning framework based on MLP-Mixer. Additionally, they introduced the multi-filter to multi-channel (MFMC) preprocessing of ECG and PPG signals. Similarly, Jia et al. (2022) estimate the blood pressure using a deep learning algorithm that learns PTT information from images of the face and palm. However, algorithms using sensor data as input are limited by their dependence on contact sensors, or the need to acquire data from multiple body parts, which restricts movement of subjects when measuring blood pressure.
Image-based blood pressure estimation methods have been conducted by combining rPPG estimation through mathematical modeling and leveraging deep learning model to estimate blood pressure from PPG signal. Schrumpf et al. (2021) conducted the study to estimate blood pressure from rPPGs by employing a rPPG extraction technique via a POS algorithm from ROI sequences and a PPG-based blood pressure estimation algorithm. Wu et al. (2022) also applied a POS algorithm to ROI sequence in 3 regions for rPPG extraction, and proposed a deep learning method to process rPPG and biometric information that can be estimated from it in parallel for blood pressure estimation. Bousefsaf et al. (2022) proposed an algorithm that extracts rPPG as G-channel averages of ROI, preprocesses them through continuous wavelet transform (CWT) in both time and frequency domains, and employs U-Net architecture to estimate continuous blood pressure signal. Chen et al. (2023) estimated blood pressure without rPPG extraction by employing spatiotemporal mapping of ROI sequences and a combined CNN and LSTM structure. While existing blood pressure estimation studies utilize mathematical modeling methods for rPPG extraction, we propose a deep learning algorithm for image-based rPPG estimation, which demonstrates improved performance in both heart rate and blood pressure estimation.
3 Proposed method
3.1 Overall training pipeline
Figure 1: Overview of the training pipeline of the proposed method.
This paper proposes a 2-stage deep learning framework which consists of DRP-Net and BBP-Net for estimating the heart rate and blood pressure from facial videos. The proposed deep learning model extracts acral and facial rPPG signals and analyzes their temporal discrepancy to estimate SBP and DBP values. Fig. 1 presents the overview of the training pipeline of the proposed method. In the preprocessing step, facial regions are detected within a video clip to define a region of interest (ROI). The DRP-Net extracts acral and facial rPPG signals from the ROI sequence, and the BBP-Net predicts SBP and DBP values. To train the DRP-Net and BBP-Net, ground truth signals are generated from ABP signals. Pseudo PPG signals are computed by detrending and bandpass filtering to supervise the training of the DRP-Net Fast Fourier transform (FFT) and power spectrum density analysis are conducted to extract HR values from pseudo PPG and rPPG signals. Additionally, peaks and valleys are detected from the ABP signals to compute the ground truth SBP and DBP values. In Fig. 1, blue and red colors indicate ground truth and predicted signals, respectively.
3.2 Preprocessing and ground truth generation
The preprocessing step extracts normalized ROI regions from the input images to remove the redundant background information. The pretrained MTCNN  (Zhang et al., 2016a) model was utilized to detect facial regions, and an ROI was decided to include all facial regions within a short video clip. The ROI regions are cropped and resized into the size of 128$\times$128, and the ROI images were normalized into the range between 0 and 1 to reduce the effect of light conditions.
On the other hand, pseudo PPG signals were generated from ABP signals by conducting detrending, bandpass filtering, and normalization. Detrending is the process of computing moving average of the ABP signals to remove high frequency noise components, and filtering was conducted to extract physiologically meaningful information from a pre-defined frequency band. Following the previous literatureHu et al. (2022), we used the smaller window size of 150 for detrending and extracted the information within the frequency band between 0.5 hz and 3 hz. The filtered signals were normalized into the range between -2 and 2 to generate pseudo PPG signals. Additionally, ground truth HR was computed by detecting the frequency value corresponding to the maximum power spectrum density. Ground truth values for SBP and DBP were computed by averaging peak and valley values within a time interval.
Figure 2: Augmentation of bradycardia and tachycardia data.
3.3 Data augmentation
Biometric information such as heart rate follows a Gaussian distribution and requires augmentation of long-tailed data due to data imbalance. In this paper, we propose upsampling and downsampling of the time axis based on frame interpolation to augment the data, and the process is shown in Fig. 2. In the training dataset, the original video sequence and signals with a window length of $T$ contain heart rate information of $r$ BPM. To augment the bradycardia data, the $\frac{T}{2}$ window length original video with a heart rate of $r$ BPM is used, and interpolation of the intermediate frames increases the sequence length to $T$ and reduces the heart rate to about $\frac{r}{2}$ BPM. The algorithm used for frame interpolation is FILM-Net (Reda et al., 2022) with transfer learning parameters. In contrast, the augmentation of tachycardia data uses the $2T$ window length original video with a heart rate of $r$ BPM, and skips an even number of frames to generate the $T$ window length sequence data with the heart rate of $2r$ BPM. The augmentation of heart rate data by upsampling and downsampling of frames has been proposed in Yu et al. (2020). The difference with our proposed method is that we use a deep learning-based frame interpolation method for upsampling.
Figure 3: Architecture of DRP-Net.
3.4 Dual remote photoplethysmography network (DRP-Net)
In this paper, we propose a dual remote photoplethysmography network (DRP-Net) to estimate phase-shifted rPPGs from image sequences, whose structure is shown in Fig. 3. It is important for the rPPG estimation algorithm to learn spatiotemporal features to capture the time-series pulse wave signal $\textbf{y}_{PPG}\in R^{T}$ in the facial image sequence $X\in R^{3\times T\times 128\times 128}$. DRP-Net is a 3 dimensional CNN model, consisting of a feature extractor, spatial and temporal attention modules, and an rPPG prediction heads. The feature extractor employs atrous convolution to learn sparse local features over a large receptive field, while reducing the spatial dimension and maintaining the temporal dimension. The feature extractor produces a spatiotemporal feature map $\textbf{m}\in R^{64\times T\times 4\times 4}$, and it consists of the 7 atrous convolutions and 3 max pooling layers. In addition, the intermediate feature map $\textbf{m}_{inter}\in R^{64\times T\times 32\times 32}$ from the feature extractor is utilized in the spatial and temporal attention modules in parallel. The spatial attention module consists of 3 dimensional convolution, max pooling, average pooling, and pointwise convolution for dimensionality reduction to obtain $\textbf{v}_{s}\in R^{1\times 1\times 4\times 4}$. The spatial attention vector $\boldsymbol{\alpha}_{s}\in[0,1]$ is computed by applying the sigmoid function to $\textbf{v}_{s}$, and is defined as follows.
$\boldsymbol{\alpha}_{s}=\frac{exp(\textbf{v}_{s})}{1+exp(\textbf{v}_{s})},$ (1)
where $\boldsymbol{\alpha}_{s}$ separate the regions that contain physiological information from the background regions, giving the skin regions higher weights. Similarly, we obtain $\textbf{v}_{t}\in R^{1\times t\times 1\times 1}$ from the temporal attention module, and the temporal attention vector $\boldsymbol{\alpha}_{t}\in[0,1]$ is defined as
$\boldsymbol{\alpha}_{t}=\frac{exp(\textbf{v}_{t})}{1+exp(\textbf{v}_{t})},$ (2)
where $\boldsymbol{\alpha}_{t}$ gives relatively low weight to motion noise at the frame level and optimizes the model to minimize its impact. The rPPG prediction head estimates phase-shifted facial rPPG $\hat{\textbf{y}}_{f}\in R^{t}$ and acral rPPG $\hat{\textbf{y}}_{a}\in R^{t}$, and the input $\boldsymbol{\beta}\in R^{64\times T\times 4\times 4}$ is defined as follows.
$\boldsymbol{\beta}=(\textbf{m}\otimes\boldsymbol{\alpha}_{t})\otimes\boldsymbol{\alpha}_{s},$ (3)
where $\otimes$ denotes elementwise multiplication, which gives the weights of temporal and spatial attention vectors $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\alpha}_{s}$ to spatiotemporal feature map m.
Figure 4: Architecture of BBP-Net.
3.5 Bounded blood pressure network (BBP-Net)
To estimate SBP and DBP from facial rPPG $\hat{\textbf{y}}_{f}$ and acral rPPG $\hat{\textbf{y}}_{a}$, we designed a multi-scale feature fusion (MSF)-based bounded blood pressure network (BBP-Net), which is shown in Fig. 4. BBP-Net takes the $6\times T$ set of physiological signals as a stack of velocity plethysmograms (VPGs) and acceleration plethysmograms (APGs), which are the first and second derivatives of the facial and acral rPPG. The structure of BBP-Net is designed with a backbone network based on MSF blocks and a parallel BP prediction heads, and scaling sigmoid function. The MSF block performs depth-wise separable (DWS) convolution with kernel sizes of the 3 and 5 in parallel. Then it performs a linear operation and a channel-wise softmax operation to compute the weight vector. Every MSF blocks contain the residual connections. The BP prediction head includes the computation process to obtain the weight vector by employing the bottleneck attention module (BAM) (Park et al., 2018) and residual connections in parallel. We modify the sigmoid function into a scaling sigmoid function to bound the estimated blood pressure to the normal range, and it is defined as follows.
$\hat{y}_{BP}=BP_{min}+\frac{BP_{max}-BP_{min}}{1+exp({-z}+\tau)},\,where\,BP\in\left\{SBP,DBP\right\},$ (4)
where $\hat{y}_{BP}$ is the estimated blood pressure, and ${BP}_{max}$ and ${BP}_{min}$ are heuristic parameters that represent the range of reasonable blood pressure. $z$ and $\tau$ are the logit estimated from the BP prediction head and temperature, respectively. The structure of BBP-Net is based on the architecture proposed by  Hu et al. (2022), which is designed to estimate SBP and DBP from PPG, and has achieved state-of-the-art (SOTA) performance on the UCI-dataset. While previous studies that take a single PPG of size $3\times T$ as input, we utilize the phase-shifted information of the rPPG and employ a scaling sigmoid function.
3.6 Loss functions
Since the purpose of DRP-Net and BBP-Net is regression to estimate the biometrics of rPPG, heart rate, and blood pressure, we designed a loss function based on the mean squared error (MSE). The loss function $L_{HR}$ of DRP-Net for heart rate estimation is denoted as
$L_{HR}=\frac{1}{n}\sum_{k=1}^{n}\left(\hat{y}_{HR,k}-y_{HR,k}\right)^{2},$ (5)
where $\hat{y}_{HR,k}$ and $y_{HR,k}$ are the predicted and ground truth heart rate, respectively, and $n$ is the number of data samples. The loss function in the frequency domain $L_{frequency}$ is defined based on the power spectrum of the physiological signal via FFT and can be defined as follows.
$L_{frequency}=\frac{1}{n}\sum_{k=1}^{n}\left(FFT(\hat{\textbf{y}}_{k})-FFT(\textbf{y}_{PPG,k})\right)^{2},$ (6)
where $\hat{\textbf{y}}_{k}$ and $\textbf{y}_{PPG,k}$ denote the rPPG predicted by DRP-Net and the PPG obtained from the sensor, respectively. In the case of using only frequency domain losses, the range of predicted signal’s amplitude is relatively small, therefore we propose $L_{peakvalley}$ as a loss function for scaling.
$L_{peakvalley}=\frac{1}{n}\sum_{k=1}^{n}\left[\left\{\phi(\hat{\textbf{y}}_{k}^{P})-\phi(\textbf{y}_{PPG,k}^{P})\right\}^{2}+\left\{\phi(\hat{\textbf{y}}_{k}^{V})-\phi(\textbf{y}_{PPG,k}^{V})\right\}^{2}\right],$ (7)
where $P$ and $V$ denote the peak and valley values of the signal, respectively, and the average of these values is denoted by $\phi$. $L_{peakvalley}$ is the time domain loss function but it is independent of the phase of the predicted signal. In addition, we define a loss function $L_{temporal}$ in the time domain to train an acral rPPG that simulates the scale and phase of the PPG. It is defined as
$L_{temporal}=\frac{1}{n}\sum_{k=1}^{n}\left(\hat{\textbf{y}}_{k}-\textbf{y}_{PPG,k}\right)^{2}.$ (8)
The total loss function to optimize the facial rPPG is $L_{facial}=\lambda_{1}L_{HR}+\lambda_{2}L_{frequency}+L_{peakvalley}$, which guides the model to distinguish pulse waves in the facial image sequence. On the other hand, acral rPPG utilizes $L_{acral}=\lambda_{1}L_{HR}+\lambda_{2}L_{frequency}+L_{peakvalley}+L_{temporal}$ to simulate the phase of the PPG. BBP-Net for blood pressure estimation employs Huber loss (Collins, 1976) to predict SBP and DBP which we define as $L_{BP}$. Additionally, acral rPPG scaled with SBP and DBP ranges is adjusted by minimizing $L_{ABP}$ to the ABP signal.
4 Experimental results
Experiments were conducted on a hardware environment including Intel Core i9-10940X CPU, 64 GB DDR4 RAM, and NVIDIA Geforce RTX 3090 Ti. Python and Pytorch were used for the implementation of the proposed algorithm. For experimental details, a time window of 150 (6 seconds), a batch size of 8, and a decay factor of 0.99 were used, and the learning rate of the Adam optimizer was set to 0.001 and 0.0001 for MMSE-HR and V4V, respectively. To evaluate the accuracy of estimated heart rate and blood pressure, we adopted the mean absolute error (MAE) and root mean squared error (RMSE). Moreover, the proposed method was evaluated using the Pearson correlation coefficient r.
4.1 Datasets
To demonstrate the heart rate and blood pressure estimation accuracy of the proposed method, we experimented on the MMSE-HR (Multimodal Spontaneous Expression-Heart Rate) (Zhang et al., 2016b) and V4V (Vision for Vitals) database (Revanur et al., 2021). The two datasets are sub-datasets of the MMSE database (BP4D+), which consists of synchronized facial image sequences and continuous blood pressure signals. As introduced in Section 3.1, the biometric information of PPG, heart rate, SBP, and DBP can be calculated by preprocessing the blood pressure signals. To ensure data diversity, tasks were performed to arouse amusement, pain, anger, and disgust. RGB video was recorded at the resolution of 1040$\times$1392 and the frame rate of 25 FPS, and blood pressure signals were collected at the sampling rate of 1000 Hz. The MMSE-HR database released at the CVPR 2016 for The Heart Rate Estimation Algorithm Challenge includes 102 samples from 17 male and 23 female subjects. The average signal lengths are 30 seconds. The V4V database released at the ICCV workshop 2021 for The 1st Vision for Vitals Challenge includes 1,358 data samples split into 724 training, 276 validation, and 358 test samples. The average signal lengths are 40 seconds, collected from 179 subjects.
Table 1: Comparison of heart rate estimation results on the MMSE-HR database.
Methods Window (sec) MAE (BPM) RMSE r
POS (De Haan & Jeanne, 2013) - 5.77 - 0.82
DeepPhys (Chen & McDuff, 2018) 30 4.72 8.68 0.82
Benefit of distraction (Nowara et al., 2021) 2 2.27 4.9 0.94
EfficientPhys-C (Liu et al., 2021) video level 3.48 7.21 0.86
CAN with synthetic data (McDuff et al., 2022) 30 2.26 3.7 -
PhysFormer (Yu et al., 2022) 6.4 2.84 5.36 0.92
Spatiotemporal feature (Jaiswal & Meenpal, 2022) 5 6.4 6.82 0.95
X-iPPGNet (Ouzar et al., 2023) 2 4.1 5.32 0.85
PhysFormer++ (Yu et al., 2023) 6.4 2.71 5.15 0.93
Ours (facial) 6 1.78 4.27 0.95
Ours (acral) 6 1.91 4.74 0.93
Figure 5: Pearson correlation graph of predicted and actual HR
Figure 6: Predicted facial(blue), acral(green) rPPG and PPG(red)
4.2 Quantitative results on MMSE-HR database
In this section, we compare the heart rate estimation performance of the proposed algorithm with previous methods using the public dataset MMSE-HR. To deal with the size of the MMSE-HR database, we employed the 5-fold cross-validation method for heart rate estimation, and the folds were split independently between subjects to demonstrate generalizability. Table. 1 presents the average heart rate estimation performance of facial and acral rPPG. The proposed method achieved MAE, RMSE, and r of 1.85, 4.51, and 0.948, respectively, and outperformed previous deep learning-based approaches. The Pearson correlation of predicted and actual heart rate based on facial rPPG per fold is shown in Fig. 5. This graphs represent a linear trend, indicating a significant similarity between the predicted and actual heart rate. Algorithms (Nowara et al. (2021), Jaiswal & Meenpal (2022), Ouzar et al. (2023)) with short input image sequences provide fast inference times but at the cost of higher error rates. Video transformer-based algorithms such as Yu et al. (2022) and Yu et al. (2023) show considerable performance in cross-dataset tests. However, they require high-performance computing equipment due to their extensive parameters and computational demands. Additionally, Fig. 6 shows the 2 predicted rPPG and their phase differences as the cardiac cycle increases. Comparing the 2 approaches in our proposed method in Table. 1, it indicates that the heart rate estimation error based on the facial rPPG is lower than the acral rPPG. These experimental findings suggest that forcing the rPPG phase leads to decreased model performance and emphasize the importance of inferring actual pulse waves from captured region.
Table 2: Comparison of blood pressure estimation results on the MMSE-HR database.
Methods Window (sec) SBP DBP
MAE (mmHg) RMSE MAE (mmHg) RMSE
NCBP (Rong & Li, 2021) - 17.52 22.43 12.13 15.23
NIBPP (Schrumpf et al., 2021) 7 13.6 - 10.3 -
BPE-Net (Chen et al., 2023) 6 12.35 16.55 9.54 12.22
Ours 6 10.19 13.01 7.09 8.86
Figure 7: Bland-Altman plot of predicted blood pressure(left: SBP, right: DBP)
The proposed BBP-Net for blood pressure estimation also outperformed the existing methods, achieving MAE and RMSE of 10.19 and 13.01 for SBP and 7.09 and 8.86 for DBP, respectively. Table. 2 presents the results of the blood pressure estimation experiment on the MMSE-HR database, and Fig. 7 shows the results of the Bland-Altman plot. The Rong & Li (2021) and Schrumpf et al. (2021) algorithms extracted rPPG using a mathematical approach and estimated blood pressure through handcrafted features and deep learning-based models, respectively. These studies are similar to our approach in extracting rPPG as an intermediate step. However, our method improves performance by considering phase-shifted information for blood pressure estimation. Chen et al. (2023) proposed a blood pressure estimation model based on a 2 dimensional spatiotemporal map, a video preprocessing technique based on the YUV color space. However, they neglect pulse waves at the acral sites of the body, resulting in higher SBP and DBP error rates compared to the proposed method. The experimental results in Table. 2 represent that the temporal information in the pulse wave is meaningful for non-contact blood pressure estimation.
4.3 Quantitative results on V4V database
Table 3: Comparison of heart rate estimation results on the V4V database.
Methods MAE (BPM) RMSE r
DeepPhys (Chen & McDuff, 2018) 10.2 13.25 0.454
PhysNet (Yu et al., 2019) 13.15 19.23 0.746
APNET (Kim et al., 2022) 4.89 7.68 0.738
Ours (facial) 3.83 9.59 0.75
Ours (acral) 4.13 10.14 0.725
In this section, we conducted experiments on V4V database to compare the heart rate estimation performance of the proposed DRP-Net with existing algorithms and the experimental results are shown in Table. 3. The experiments are conducted on the training and test data, and the graph 6 in Fig. 5 shows a Pearson correlation plot of the predicted and actual heart rates. The metrics for DRP-Net heart rate estimation performance using V4V are the same as for the MMSE-HR database. The proposed method using face rPPG outperformed in MAE and r, with 3.83 and 0.75, respectively. Additionally, the proposed method with acral rPPG achieved MAE, RMSE, and r of 4.13, 10.14, and 0.72, respectively This result is a slightly higher error than facial rPPG, similar to the results on the MMSE-HR database. The performance comparison between Yu et al. (2019) and DRP-Net shows that the wider receptive fields through 3D atrous convolution is beneficial for heart rate estimation than the encoder-decoder structure based on 3D CNN. The performance comparison with Kim et al. (2022) shows contrasting MAE and RMSE results. It indicates that the outliers of heart rate predicted by the proposed method are larger, while the average error is smaller.
Table 4: Comparison of blood pressure estimation results on the V4V database.
Methods SBP DBP
MAE (mmHg) RMSE MAE (mmHg) RMSE
AlexNet 18.582 - 13.46 -
ResNet 15.456 - 12.528 -
LSTM 14.579 - 13.119 -
NIBPP (Schrumpf et al., 2021) 31.359 - 20.621 -
Hamoud et al. (2023)
(Combination of model 1,3)
15.121 - 11.169 -
Hamoud et al. (2023)
13.749 - 14.833 -
Ours 13.640 16.782 9.402 11.903
The blood pressure estimation results of BBP-Net is shown in Table. 4. Our proposed algorithm achieved MAE and RMSE of 13.64, 16.78 and 9.40, 11.90 for SBP and DBP, respectively. The performance of the baseline algorithms (AlexNet, ResNet, LSTM) is reported in Hamoud et al. (2023). Blood pressure estimation performance of BBP-Net outperforms baseline algorithms such as AlexNet, ResNet, and LSTM, as well as Schrumpf et al. (2021) and the algorithm proposed by Hamoud et al. (2023). While the SBP and DBP estimation performance presented in the Hamoud et al. (2023) has the limitation that each performance is a combination of different models, the BBP-Net estimates both blood pressures simultaneously. As shown in the experimental results on the V4V database, the rPPG estimated from the 2 body parts and its phase-shifted information are effective in improving the estimation performance of blood pressure.
4.4 Ablation study
Table 5: Comparative results of mean absolute error in heart rate estimation task across various window sizes and folds.
Window
rPPG Fold 0 Fold 1 Fold 2 Fold 3 Fold 4 Average
2 facial 11.22 13.71 9.79 11.96 15.15 12.37
acral 12.17 11.31 10.34 11.73 14.10 11.93
4 facial 1.86 2.56 1.90 2.32 4.26 2.58
acral 1.79 2.79 1.95 2.17 4.05 2.55
6 facial 1.58 1.08 0.88 1.79 3.58 1.78
acral 1.46 1.00 0.82 1.97 4.29 1.91
8 facial 1.20 1.15 0.91 1.06 3.11 1.49
acral 1.13 1.04 0.58 1.45 3.20 1.48
Table 6: Comparison of MAE in heart rate estimation with and without data augmentation.
Augmentation rPPG MAE (BPM)
X facial 2.149
acral 2.002
O facial 1.783
acral 1.909
Table 7: Comparison of MAE in heart rate estimation with and without peak valley loss.
Peak valley loss rPPG MAE (BPM)
X facial 3.670
acral 4.178
O facial 1.783
acral 1.909
The window size of the input image sequence is an important heuristic parameter in terms of the computational power and the latency of model. Smaller window sizes have the advantage of lower latency in inference time, while the loss of temporal information can decrease the performance. Table. 5 presents the performance with different window size and fold, and it demonstrates that the average heart rate estimation performance increases as the window size increases. The reduced performance observed at a window size of 2 seconds is due to the dilation ratio of DRP-Net being relatively large compared to the sequence information. However, estimation of blood pressure requires long temporal information of rPPG, therefore we selected window size of 6 seconds and a dilation ratio of 4. Table. 6 indicates the results of heart rate estimation of DRP-Net with and without data augmentation techniques. The performance of all estimated rPPG improved after data augmentation. This result demonstrate that training with data from bradycardia and tachycardia via frame interpolation and down-sampling improves the generalization performance of the model. In addition, the experimental results of excluding the peak and valley guides from the loss function are demonstrated in Table. 7 to investigate the impact of $L_{peakvalley}$. The average MAE without $L_{peakvalley}$ increases about 2 BPM, which indicates that the amplitude information in the PPG contributes the improvement of heart rate estimation performance.
Table 8: Comparative results of MAE in blood pressure estimation for the input rPPG types.
Input rPPG SBP MAE (mmHg) DBP MAE (mmHg)
facial rPPG 12.323 9.503
acral rPPG 13.190 8.669
both 10.192 7.090
Table 9: Comparison of blood pressure estimation performance with and without the scaling sigmoid function.
Scaling sigmoid function SBP DBP
MAE (mmHg) RMSE MAE (mmHg) RMSE
X 19.57 25.34 10.15 13.34
O 10.19 13.01 7.09 8.86
The ablation study for the blood pressure estimation model consists of the types of input rPPG and the existence of the scaling sigmoid function. Table. 8 presents the estimation error of the SBP and DBP based on the facial or acral rPPG, and the estimation performance of both cases are lower than using dual rPPG. This results indicate that the phase-shifted information of pulse waves in different body parts contributes to improved estimation performance of SBP and DBP. Additionally, Table. 9 shows the performance of blood pressure estimation with and without the scaling sigmoid function. Without the scaling sigmoid function, the MAE for SBP and DBP was significantly increased to 9 mmHg and 3 mmHg, respectively. These experimental results suggest the importance of guiding the model towards a reasonable blood pressure range.
4.5 International standards and blood pressure estimation performance
Table 10: Comparison results of blood pressure estimation performance with BHS standards.
$\leq$ 5 mmHg $\leq$ 10 mmHg $\leq$ 15 mmHg
MMSE-HR SBP 46.02 % 65.49 % 75.22 %
DBP 81.42 % 92.04 % 94.69 %
V4V SBP 65.24 % 72.94 % 81.39 %
DBP 64.17 % 80.11 % 86.63 %
BHS standard grade A 60 % 85 % 95 %
grade B 50 % 75 % 90 %
grade C 40 % 65 % 85 %
Table 11: Comparison results of blood pressure estimation performance with AAMI standards.
ME (mmHg) SDE (mmHg)
MMSE-HR SBP -0.13 16.78
DBP 1.58 11.8
V4V SBP 5.96 11.57
DBP -2.4 8.53
AAMI standard $\leq$5 $\leq$8
We compare the performance of our proposed blood pressure estimation algorithm with international standards, the British Hypertension Society (BHS)  (O’Brien et al., 1990) and The Association for the Advancement of Medical Instrumentation (AAMI)  (Stergiou et al., 2018). The BHS standard evaluates the absolute number of blood pressure estimation errors as a cumulative percentage, and Table. 10 presents the standards and blood pressure estimation performance. On the MMSE-HR database, the proposed method achieves grade B with 81 % DBP estimation performance for errors of 5 mmHg or less. Additionally, DBP estimation performance on the V4V database also reached grade C. The AAMI standard in Table. 11 indicates that blood pressure estimation performance meets the standard when the mean error (ME) and standard deviation error (SDE) are under 5 and 8 mmHg, respectively. In the MMSE-HR database, the estimation performance of SBP and DBP for ME is -0.13 mmHg and 1.58 mmHg, respectively, which meets the standard. The result on the V4V database shows that ME is 5.96 mmHg at SBP and -2.4 mmHg at DBP.
5 Conclusion
This paper proposed the deep learning framework based on facial images for contactless heart rate and blood pressure estimation. We assumed that phase differences exist between pulse waves due to inconsistencies in data measurement sites. Based on this assumption, we improved blood pressure estimation performance and demonstrated with extensive experiments. To estimate pulse waves of 2 physiological sites from facial images, we proposed the DRP-Net that focuses on spatiotemporal feature extraction. DRP-Net estimates facial rPPG using information in the time-global frequency domain to predict pulse waves from the face, and predicts acral rPPG that mimic the phase of PPG. As a result, the proposed model achieved the MAE of 1.78 and 1.91 BPM based on acral and facial rPPG in heart rate estimation task. Moreover, we proposed BBP-Net to estimate blood pressure based on the phase analysis and temporal features of the dual rPPG, applying the modified sigmoid function to bound the predicted blood pressure to the reasonable range. In the blood pressure estimation task, we achieved MAE of SBP and DBP of 10.19 mmHg and 7.09 mmHg, respectively. However, the public datasets used in this study have the limitation in that they do not include pulse wave data from the face, making it unable to verify the accuracy of the predicted facial rPPG phase. We were also constrained by the absence of other public datasets containing both facial images and blood pressure data, limiting experimentation in various domains. Inspired by these limitations, we have plan to collect real-world dataset which includes facial videos, PPG, and blood pressure with various data collection scenarios and develop sensor acquisition system in our future research. Additionally, we will conduct research on improving the performance of physiological measurement system using the collected real-world dataset.

### 06 Mar 2024 v2

@@
 Human health can be critically affected by cardiovascular diseases, such as hypertension, arrhythmias, and stroke.
-Heart rate and blood pressure are important biometric information for the monitoring of cardiovascular system and early diagnosis of cardiovascular diseases.
-Existing methods for estimating the heart rate are based on electrocardiography and photoplethyomography, which require contacting the sensor to the skin surface.
-Moreover, catheter and cuff-based methods for measuring blood pressure cause inconvenience and have limited applicability.
-Therefore, in this thesis, we propose a vision-based method for estimating the heart rate and blood pressure.
-This thesis proposes a 2-stage deep learning framework consisting of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net).
-In the first stage, DRP-Net infers remote photoplethysmography (rPPG) signals for the acral and facial regions, and these phase-shifted rPPG signals are utilized to estimate the heart rate.
-In the second stage, BBP-Net integrates temporal features and analyzes phase discrepancy between the acral and facial rPPG signals to estimate SBP and DBP values.
-To improve the accuracy of estimating the heart rate, we employed a data augmentation method based on a frame interpolation model.
-Moreover, we designed BBP-Net to infer blood pressure within a predefined range by incorporating a scaled sigmoid function.
-To demonstrate the effectiveness of the proposed method, experiments were conducted on the MMSE-HR and V4V datasets.
-Our method resulted in estimating the heart rate with the mean absolute error (MAE) of 1.78 BPM, reducing the MAE by 34.31 % compared to the recent method, on the MMSE-HR dataset.
-The MAE for estimating the systolic blood pressure (SBP) and diastolic blood pressure (DBP) were 10.19 mmHg and 7.09 mmHg.
-On the V4V dataset, the MAE for the heart rate, SBP, and DBP were 3.83 BPM, 13.64 mmHg, and 9.4 mmHg, respectively.
-We expect that the proposed method will contribute to the advancement of sensing technology for the monitoring of human health.
+Heart rate (HR) and blood pressure (BP) are important physiological information for monitoring of cardiovascular system and early diagnosis of cardiovascular diseases.
+Previous methods for estimating heart rate are based on electrocardiography and photoplethysmography, which require contacting sensors to skin surface.
+Existing catheter and cuff-based methods for measuring blood pressure cause inconvenience and are difficult to be utilized in daily life.
+To address these limitations, this paper proposes a 2-stage deep learning framework for estimating heart rate and blood pressure from facial video.
+The proposed algorithm consists of a dual remote photoplethysmography network (DRP-Net) and bounded blood pressure network (BBP-Net).
+DRP-Net infers remote photoplethysmography (rPPG) signals at acral and facial regions, and these phase-shifted rPPG signals are utilized to estimate heart rate.
+BBP-Net integrates temporal features and analyzes phase discrepancy between the acral and facial rPPG signals to estimate systolic blood pressure (SBP) and diastolic blood pressure (DBP).
+We augmented facial video datasets in temporal aspect by utilizing a frame interpolation model to improve the accuracy of estimating heart rate.
+Moreover, we reduced the error in estimating blood pressure by incorporating a scaled sigmoid function in the BBP-Net.
+Experiments were conducted on MMSE-HR and V4V datasets to demonstrate the effectiveness of the proposed method.
+Our method achieved the state-of-the-art performance for estimating heart rate and blood pressure with significant margins compared to previous methods.
+Code is available at https://github.com/GyutaeHwang/phase_shifted_rPPG.
 keywords:
@@
 Moreover, cardiovascular diseases associated with blood vessels, such as hypertension and stroke, are contribute to increasing the global mortality rate (Gaidai et al., 2023).
-However, these diseases can be detected at an early stage through the monitoring of the heart rate and blood pressure using healthcare devices such as smartwatches.
+However, these diseases can be detected at an early stage through the monitoring of the heart rate (HR) and blood pressure (BP) using healthcare devices such as smartwatches.
 For example, arrhythmias caused by abnormalities in the sinoatrial node can be prevented through the measurement of bradycardia, tachyarrhythmia, and irregular heart rates.
 Additionally, consistent readings of systolic blood pressure (SBP) above 120 mmHg and diastolic blood pressure (DBP) above 90 mmHg may indicate the potential presence of hypertension.
-Recently, interest in cardiovascular diseases has led to growing attention on research for healthcare services (Etemadi et al., 2023; Aggarwal & Pandey, 2023) and estimation of biometric information (Alnaggar et al., 2023; Gupta et al., 2023).
-Heart rate and blood pressure are main biometric information for the monitoring of cardiovascular diseases (Perret-Guillaume et al., 2009; Fuchs & Whelton, 2020), and these information can be measured by utilizing various physiological sensors and medical equipment.
-Heart rate is commonly measured by using electrocardiography (ECG) and photoplethysmography (PPG).
-ECG uses electrodes attached to the body to record the electrical activity caused by contractions and relaxations of the heart.
-PPG is a non-invasive method that uses a light source and photodetector at the skin surface to measure the volumetric variations of blood in micro vessels.
-The heart rate is measured by computing peak-to-peak intervals in the time domain or by analyzing the power spectral density in the frequency domain from the ECG or PPG signals.
-On the other hand, blood pressure can be measured based on an oscillometric method, which records the magnitude of pressure oscillations using a cuff, or an invasive catheter-based approach.
-The catheter-based method involves the direct insertion of a sensor into an artery to measure continuous arterial blood pressure (ABP).
-Peak and valley values of the ABP signal are utilized as SBP and DBP, respectively.
-The blood pressure can also be estimated from pulse transit time (PTT) measured by ECG or PPG sensors at the heart and acral sites.
-PTT is the time taking for a blood pulse wave to travel from the heart to an acral site, and it known that PTT is closely correlated with blood pressure (Geddes et al., 1981).
-Although there have been proposed BP estimation methods by analyzing PTT from ECG or PPG signals (Barvik et al., 2021; Ganti et al., 2020), these approaches have the intrinsic limitation of requiring skin contact.
-Recently, camera sensors have been utilized to obtain physiological information, and it is called remote photoplethysmography (rPPG).
-From a facial video, the rPPG technique extracts subtle variations in skin color induced by cardiovascular pulse.
+Recently, interest in cardiovascular diseases has led to growing attention on research for healthcare services (Etemadi et al., 2023; Aggarwal & Pandey, 2023) and estimation of physiological information (Alnaggar et al., 2023; Gupta et al., 2023).
+Heart rate and blood pressure are main physiological information for the monitoring of cardiovascular diseases (Perret-Guillaume et al., 2009; Fuchs & Whelton, 2020), and these information can be measured by utilizing various physiological sensors and medical equipment.
+Heart rate is commonly measured by utilizing electrocardiography (ECG) and photoplethysmography (PPG).
+ECG employs electrodes attached to the body to record the electrical activity caused by contractions and relaxations of the heart.
+PPG is a non-invasive method that uses a light source and photodetector attached at skin surfaces to measure the volumetric variations of blood in micro vessels.
+The heart rate can be measured by computing peak-to-peak intervals in the time domain or by analyzing power spectrum in the frequency domain from ECG or PPG signals.
+On the other hand, blood pressure can be measured based on an oscillometric method, which records the magnitude of oscillations using a blood pressure cuff.
+A catheter-based method is an invasive approach to measure blood pressure, and it involves direct insertion of a sensor into an artery to measure real-time arterial blood pressure (ABP).
+SBP and DBP can be computed from peak and valley values of the ABP signals.
+Pulse transit time (PTT) is temporal delay of a blood pulse wave to travel from the heart to an acral site such as a fingertip, and it is known that PTT is closely correlated with blood pressure (Geddes et al., 1981).
+Although there have been proposed blood pressure estimation methods by analyzing PTT from ECG or PPG signals (Barvik et al., 2021; Ganti et al., 2020), these approaches have the intrinsic limitation of requiring skin contacts.
+Recently, camera sensors have been utilized to obtain physiological signals, and it is called remote photoplethysmography (rPPG).
+From a facial video, the rPPG technique extracts subtle variations in skin color induced by cardiac pulse.
 The extracted temporal changes in pixel intensities are then transformed into continuous waveforms analogous to conventional PPG signals.
 The precision of the rPPG technique can be critically influenced by many factors such as ambient light conditions, motion artifacts, and different skin tones.
-Despite these challenges, rPPG receives much attention due to the non-contact and non-invasive attributes that facilitate remote monitoring of biometric information.
+Despite these challenges, rPPG receives much attention due to the non-contact and non-invasive attributes that facilitate remote monitoring of physiological information.
 Deep learning methods have been utilized to extract rPPG signals and to estimate the heart rate (Chen & McDuff, 2018; Yu et al., 2022; Ouzar et al., 2023).
@@
 PPG signals in the existing datasets were collected at acral sites such as fingertips, and therefore, there exist temporal discrepancies between the pulse waves in the PPG signals and facial videos.
-However, most previous methods ignore the time shift between the two modalities and adopt a time domain loss between PPG signals measured at an acral site and rPPG signals extracted from facial videos.
+However, most previous methods neglect the time shift between the two modalities and adopt a time domain loss between PPG signals measured at an acral site and rPPG signals extracted from facial videos.
 For example, Lu et al.
@@
 In this paper, we introduced two types of phase-shifted rPPG signals, referred to as acral rPPG and facial rPPG, to consider the temporal discrepancy between facial videos and PPG signals measured at an acral site.
-The acral rPPG is guided by a time domain loss function, while the facial rPPG is guided by frequency domain features which have temporally global characteristics.
+The acral rPPG signal is guided by a time domain loss function, while the facial rPPG signal is guided by frequency domain features which have temporally global characteristics.
 On the other hand, blood pressure can be estimated from PPG signals, both ECG and PPG signals, or facial videos.
 While PPG-based methods are widely utilized in wearable devices, this approach has the limitation of requiring physical contact to obtain PPG signals.
-Although the approach of utilizing both ECG and PPG signals can measure blood pressure based on PTT information, it also has the disadvantage of requiring contact with ECG and PPG sensors.
+Although the approach of utilizing both ECG and PPG signals can measure blood pressure based on PTT, it also has the disadvantage of requiring contact with ECG and PPG sensors.
 Recently, camera-based methods have been proposed to estimate blood pressure in a non-contact manner by utilizing spatiotemporal features in facial videos.
@@
 In the first stage, DRP-Net infers the acral and facial rPPG signals, and it is used to estimate the heart rate.
-In the second stage, BBP-Net integrates temporal features based on a multi-scale fusion (MSF) module and analyzes phase discrepancy between the acral and facial rPPG signals to estimate SBP and DBP values.
+In the second stage, BBP-Net integrates temporal features based on a multi-scale fusion (MSF) module and analyzes phase discrepancy between the acral and facial rPPG signals to estimate SBP and DBP.
 To improve the precision of blood pressure estimation, a scaled sigmoid function is employed in the BBP-Net to constrain the estimate into a certain range.
@@
   • 2.
-    We employ a frame interpolation algorithm for the augmentation of bradycardia and pyknocardia video clips.
+    We employ a frame interpolation algorithm for the augmentation of bradycardia and tachycardia video clips.
   • 3.
-    We introduce the concept of phase-shifted rPPG for considering the temporal discrepancy between pulse waves in facial videos and PPG signals.
+    We introduce the concept of phase-shifted rPPG signals for considering the temporal discrepancy between pulse waves in facial videos and PPG signals.
   • 4.
@@
 2.1 Estimation of rPPG signals from facial videos
-Image processing techniques and mathematical models have been utilized to analyze subtle variations in color intensities and to estimate rPPG signals.
+The conventional rPPG extraction method analyzed subtle variations in color intensities using image processing techniques and mathematical modeling.
 Poh et al.
-(2010) computed pulse waves by averaging on RGB color channels, and independent component analysis was conducted to extract rPPG signals.
-Lewandowska et al.
-(2011) estimated rPPG signals by conducting a channel selection process and principal component analysis on color intensities in a forehead region.
-De Haan & Jeanne (2013) proposed a chrominance-based method, called CHROM, to improve the performance of rPPG against motion artifacts.
+(2010) computed pulse waves by averaging RGB color channels, and independent component analysis was conducted to extract rPPG signals.
+On the other hand, Lewandowska et al.
+(2011) estimated rPPG signals by conducting a channel selection process and principal component analysis on color intensities in the forehead region.
+De Haan & Jeanne (2013) proposed a chrominance-based method, called CHROM, to improve the performance of extracting rPPG signals against motion artifacts.
 Wang et al.
 (2016) proposed the POS algorithm, which analyzes the plane orthogonal to the skin tone in the normalized RGB space.
-Meanwhile, with the success of deep learning methods for computer vision tasks, data driven approaches started to be applied in the estimation of rPPG signals.
-The deep learning-based rPPG estimation technique focuses on learning trends in PPG that are robust to motion and external brightness changes from video sequence data with spatiotemporal characteristics.
-Chen & McDuff (2018) introduced a convolutional attention network (CAN) that employs a Siamese-structured 2 dimensional CNN model.
-This model takes both a difference map of adjacent viewpoints and the original viewpoint image.
+Meanwhile, the success of deep learning methods for computer vision tasks, data driven approaches began to be applied in the estimation of rPPG signals.
+The deep learning-based rPPG estimation technique focuses on learning trends of PPG signals that are strongly reactive to motion and external brightness changes from video sequence data with spatiotemporal characteristics.
+Chen & McDuff (2018) introduced the convolutional attention network (CAN) that employs a Siamese-structured two-dimensional convolutional neural network (CNN) model.
+This model takes both the original temporal point image and a difference map of adjacent points.
 Spatial attentions are utilized to emphasize skin regions in the difference images, effectively integrating the models.
@@
 (2021) aimed to estimate spatial noise factors including motion and illumination using the inverse of spatial attentions of the CAN model.
-They integrated the LSTM to enhance spatial denoising and improve temporal robustness in PPG.
+They integrated the LSTM to enhance spatial denoising and improve temporal robustness in PPG signals.
 Yu et al.
-(2019) proposed PhysNet, which employs a 3D CNN-based encoder to extract spatiotemporal features and a decoder along the time axis to estimate rPPG.
+(2019) proposed PhysNet3D, which employs a three-dimensional CNN-based encoder to extract spatiotemporal features and a deconvolution of the time axis to estimate rPPG signals.
 They also introduced a loss function based on the Pearson correlation coefficient.
 Yu et al.
-(2022) proposed a loss function based on temporal difference convolution (TDC) using neighboring frame differences of a 4 dimensional feature map as query and key, and KL divergence of periodic component distributions.
+(2022) proposed a video transformer based on temporal difference convolution (TDC), which utilizes adjacent frame differences of a 4-dimensional feature map as query and key.
+They also proposed a KL divergence-based loss function for periodic component distributions.
 Yu et al.
-(2023) is a SlowFast Network for temporal improvement of the estimated signal, which has shown promising heart rate estimation accuracy even in cross-domain.
-However, existing studies primarily focuses on providing performance results of heart rate estimation using the power spectrum of rPPG, and less consideration is given to the blood pressure estimation steps based on rPPG.
+(2023) utilized a SlowFast Network to improve the temporal representation of the estimated signal, which has demonstrated promising accuracy in heart rate estimation, even in cross-domain experiments.
+However, existing studies primarily focus on providing performance results of heart rate estimation using the power spectrum of rPPG signals.
+Therefore, scholars have given less consideration to the blood pressure estimation steps based on rPPG signals.
 2.2 Deep learning methods for estimating blood pressure
-For non-invasive blood pressure estimation, researchers have attempted to develop deep learning algorithms utilizing physiological signal measured from various sensors.
+Non-invasive blood pressure estimation methods aim to develop deep learning algorithms utilizing physiological signals measured from user-friendly sensors.
 Miao et al.
-(2020) proposed a deep learning model based on ResNet and LSTM for continuous blood pressure estimation from single channel ECG.
+(2020) proposed a deep learning model based on ResNet and Long Short-Term Memory (LSTM) for continuous blood pressure estimation from single channel ECG signal.
 Panwar et al.
-(2020) introduced PP-Net, which consists of a 1D convolution and LSTM, to estimate SBP, DBP, and heart rate from PPG signals.
+(2020) introduced PP-Net, which consists of a one-dimensional convolution blocks and LSTM, to estimate SBP, DBP, and heart rate from PPG signals.
 Huang et al.
-(2022) utilized PTT information for blood pressure estimation and proposed a deep learning framework based on MLP-Mixer.
-Additionally, they introduced the multi-filter to multi-channel (MFMC) preprocessing of ECG and PPG signals.
+(2022) utilized the concept of PTT for blood pressure estimation and proposed a deep learning framework based on MLP-Mixer.
+Additionally, the researchers introduced the multi-filter to multi-channel (MFMC) technique for preprocessing ECG and PPG signals.
 Similarly, Jia et al.
-(2022) estimate the blood pressure using a deep learning algorithm that learns PTT information from images of the face and palm.
-However, algorithms using sensor data as input are limited by their dependence on contact sensors, or the need to acquire data from multiple body parts, which restricts movement of subjects when measuring blood pressure.
-Image-based blood pressure estimation methods have been conducted by combining rPPG estimation through mathematical modeling and leveraging deep learning model to estimate blood pressure from PPG signal.
+(2022) estimate blood pressure using a proposed deep learning model that analyzes PTT from image sequences of the face and palm.
+However, the mentioned models require data to be acquired from contact sensors or multiple physiological sites, which have limitations that restrict the user’s movements.
+Image-based blood pressure estimation methods have been developed by combining conventional rPPG extraction methods with deep learning models that estimate blood pressure from PPG signals.
 Schrumpf et al.
-(2021) conducted the study to estimate blood pressure from rPPGs by employing a rPPG extraction technique via a POS algorithm from ROI sequences and a PPG-based blood pressure estimation algorithm.
+(2021) conducted a study to estimate blood pressure from rPPG signals by employing a rPPG extraction technique via a POS algorithm from ROI sequences and a PPG-based blood pressure estimation model.
 Wu et al.
-(2022) also applied a POS algorithm to ROI sequence in 3 regions for rPPG extraction, and proposed a deep learning method to process rPPG and biometric information that can be estimated from it in parallel for blood pressure estimation.
+(2022) also applied a POS algorithm to three ROI sequences for rPPG extraction and proposed FS-Net, which processes rPPG signals and physiological information in parallel for blood pressure estimation.
 Bousefsaf et al.
-(2022) proposed an algorithm that extracts rPPG as G-channel averages of ROI, preprocesses them through continuous wavelet transform (CWT) in both time and frequency domains, and employs U-Net architecture to estimate continuous blood pressure signal.
+(2022) used green channel average for rPPG extraction and proposed a method to restore continuous ABP signal based on continuous wavelet transform and U-Net.
 Chen et al.
 (2023) estimated blood pressure without rPPG extraction by employing spatiotemporal mapping of ROI sequences and a combined CNN and LSTM structure.
-While existing blood pressure estimation studies utilize mathematical modeling methods for rPPG extraction, we propose a deep learning algorithm for image-based rPPG estimation, which demonstrates improved performance in both heart rate and blood pressure estimation.
+While existing blood pressure estimation studies utilize non-parametric methods for rPPG extraction, we propose a deep learning algorithm for image-based rPPG estimation, which demonstrates improved performance in both heart rate and blood pressure estimation.
 3 Proposed method
 3.1 Overall training pipeline
-Figure 1: Overview of the training pipeline of the proposed method.
-This paper proposes a 2-stage deep learning framework which consists of DRP-Net and BBP-Net for estimating the heart rate and blood pressure from facial videos.
-The proposed deep learning model extracts acral and facial rPPG signals and analyzes their temporal discrepancy to estimate SBP and DBP values.
+This paper proposes a two-stage deep learning framework consisting of DRP-Net and BBP-Net to estimate heart rate and blood pressure from facial videos.
+The proposed deep learning model extracts acral and facial rPPG signals and analyzes their temporal discrepancy to estimate SBP and DBP.
 Fig.
@@
 In the preprocessing step, facial regions are detected within a video clip to define a region of interest (ROI).
-The DRP-Net extracts acral and facial rPPG signals from the ROI sequence, and the BBP-Net predicts SBP and DBP values.
-To train the DRP-Net and BBP-Net, ground truth signals are generated from ABP signals.
-Pseudo PPG signals are computed by detrending and bandpass filtering to supervise the training of the DRP-Net Fast Fourier transform (FFT) and power spectrum density analysis are conducted to extract HR values from pseudo PPG and rPPG signals.
-Additionally, peaks and valleys are detected from the ABP signals to compute the ground truth SBP and DBP values.
-In Fig.
-1, blue and red colors indicate ground truth and predicted signals, respectively.
+The DRP-Net learns spatiotemporal features from the ROI sequence and extracts acral and facial rPPG signals via dual prediction heads.
+Afterwards, The BBP-Net learns temporal multi-scale information from phase-shifted rPPG signals to predict SBP and DBP.
+To generate ground truth data, ABP signals are utilized to calculate pseudo PPG signals, heart rate, SBP and DBP.
+Figure 1: Overview of the training pipeline of the proposed method. Blue and red colors indicate ground truth and predicted physiological information, respectively.
 3.2 Preprocessing and ground truth generation
@@
 The ROI regions are cropped and resized into the size of 128$\times$128, and the ROI images were normalized into the range between 0 and 1 to reduce the effect of light conditions.
-On the other hand, pseudo PPG signals were generated from ABP signals by conducting detrending, bandpass filtering, and normalization.
-Detrending is the process of computing moving average of the ABP signals to remove high frequency noise components, and filtering was conducted to extract physiologically meaningful information from a pre-defined frequency band.
-Following the previous literatureHu et al.
-(2022), we used the smaller window size of 150 for detrending and extracted the information within the frequency band between 0.5 hz and 3 hz.
-The filtered signals were normalized into the range between -2 and 2 to generate pseudo PPG signals.
-Additionally, ground truth HR was computed by detecting the frequency value corresponding to the maximum power spectrum density.
-Ground truth values for SBP and DBP were computed by averaging peak and valley values within a time interval.
+The sequence of ROI images was sampled at 25 frames per second (FPS) and split into window length of 150 frames (6 seconds) following the previous method (Chen et al., 2023).
+The datasets utilized in this research contain ABP signals which carry various physiological information.
+The ABP signal was sampled and split at the same rate as the FPS and window length of 6 seconds of a facial video.
+Pseudo PPG signals were generated by conducting detrending, bandpass filtering, and standardization of ABP signals.
+Detrending is the process of computing time-varying finite-impulse response of the pseudo PPG signals  (Tarvainen et al., 2002) to remove high frequency noise components.
+Filtering was conducted within a pre-defined frequency band between 0.5 Hz and 3.0 Hz, enabling the acquisition of a wider range of physiologically meaningful information compared to the previous method (Nowara et al., 2021).
+The Fast Fourier Transform is used to extract the heart rate, which is computed by detecting the frequency corresponding to the maximum amplitude of the power spectral density (PSD) from the pseudo PPG signal.
+Ground truth values for SBP and DBP were computed by averaging the detected peaks and valleys within a window.
 Figure 2: Augmentation of bradycardia and tachycardia data.
 3.3 Data augmentation
-Biometric information such as heart rate follows a Gaussian distribution and requires augmentation of long-tailed data due to data imbalance.
-In this paper, we propose upsampling and downsampling of the time axis based on frame interpolation to augment the data, and the process is shown in Fig.
+We propose temporal upsampling and downsampling based on frame interpolation as an augmentation method for bradycardia and tachycardia data, and the process is shown in Fig.
 2.
-In the training dataset, the original video sequence and signals with a window length of $T$ contain heart rate information of $r$ BPM.
-To augment the bradycardia data, the $\frac{T}{2}$ window length original video with a heart rate of $r$ BPM is used, and interpolation of the intermediate frames increases the sequence length to $T$ and reduces the heart rate to about $\frac{r}{2}$ BPM.
-The algorithm used for frame interpolation is FILM-Net (Reda et al., 2022) with transfer learning parameters.
-In contrast, the augmentation of tachycardia data uses the $2T$ window length original video with a heart rate of $r$ BPM, and skips an even number of frames to generate the $T$ window length sequence data with the heart rate of $2r$ BPM.
-The augmentation of heart rate data by upsampling and downsampling of frames has been proposed in Yu et al.
-(2020).
-The difference with our proposed method is that we use a deep learning-based frame interpolation method for upsampling.
+Generally, the normal heart rate ranges between 60 and 100 BPM (Guyton, 2006), and bradycardia and tachycardia refer to heart rates lower and faster than normal heart rate, respectively.
+In the training dataset, An ROI sequence of the length of $T$ corresponding to the heart rate $r$ is augmented to generate bradycardia and tachycardia data which correspond to the heart rate of $\frac{T}{2}$ BPM and $2r$ BPM.
+To augment the bradycardia data, the $\frac{T}{2}$ frames of original ROI sequence is used, and expand the length of the window to $T$ frames via a frame interpolation method.
+The duration of the cardiac cycle decreases inversely with the increase in the sampling rate of the ROI sequence.
+The algorithm used for frame interpolation is pretrained FILM-Net (Reda et al., 2022).
+In contrast, the augmentation of tachycardia data downsamples $2T$ frames of the original ROI sequence by a factor of 2.
+The augmentation methods of remote physiological measurement include frame sampling (Yu et al., 2020; Lokendra & Puneet, 2022), color jitter (Yue et al., 2023), ROI masking (Chen et al., 2023), and synthetic data (Perche et al., 2021).
+The datasets utilized in this paper mostly contain a normal range of heart rates.
+To augment bradycardia and tachycardia data, we have improved the frame sampling methods.
 Figure 3: Architecture of DRP-Net.
 3.4 Dual remote photoplethysmography network (DRP-Net)
-In this paper, we propose a dual remote photoplethysmography network (DRP-Net) to estimate phase-shifted rPPGs from image sequences, whose structure is shown in Fig.
+In this paper, we propose a dual remote photoplethysmography network (DRP-Net) to estimate phase-shifted rPPG signals from image sequences, whose structure is shown in Fig.
 3.
-It is important for the rPPG estimation algorithm to learn spatiotemporal features to capture the time-series pulse wave signal $\textbf{y}_{PPG}\in R^{T}$ in the facial image sequence $X\in R^{3\times T\times 128\times 128}$.
-DRP-Net is a 3 dimensional CNN model, consisting of a feature extractor, spatial and temporal attention modules, and an rPPG prediction heads.
+It is important for the rPPG estimation model to learn spatiotemporal features to capture the time-series pulse wave signal in the facial image sequence $\textbf{X}\in\mathbb{R}^{3\times T\times H\times W}$.
+$H$ and $W$ denote height and width, and we set both of these hyperparameters of 128 in our experiments.
+DRP-Net is a three-dimensional CNN model, consisting of a feature extractor, spatial and temporal attention modules, and rPPG prediction heads.
 The feature extractor employs atrous convolution to learn sparse local features over a large receptive field, while reducing the spatial dimension and maintaining the temporal dimension.
-The feature extractor produces a spatiotemporal feature map $\textbf{m}\in R^{64\times T\times 4\times 4}$, and it consists of the 7 atrous convolutions and 3 max pooling layers.
-In addition, the intermediate feature map $\textbf{m}_{inter}\in R^{64\times T\times 32\times 32}$ from the feature extractor is utilized in the spatial and temporal attention modules in parallel.
-The spatial attention module consists of 3 dimensional convolution, max pooling, average pooling, and pointwise convolution for dimensionality reduction to obtain $\textbf{v}_{s}\in R^{1\times 1\times 4\times 4}$.
-The spatial attention vector $\boldsymbol{\alpha}_{s}\in[0,1]$ is computed by applying the sigmoid function to $\textbf{v}_{s}$, and is defined as follows.
-$\boldsymbol{\alpha}_{s}=\frac{exp(\textbf{v}_{s})}{1+exp(\textbf{v}_{s})},$ (1)
-where $\boldsymbol{\alpha}_{s}$ separate the regions that contain physiological information from the background regions, giving the skin regions higher weights.
-Similarly, we obtain $\textbf{v}_{t}\in R^{1\times t\times 1\times 1}$ from the temporal attention module, and the temporal attention vector $\boldsymbol{\alpha}_{t}\in[0,1]$ is defined as
-$\boldsymbol{\alpha}_{t}=\frac{exp(\textbf{v}_{t})}{1+exp(\textbf{v}_{t})},$ (2)
-where $\boldsymbol{\alpha}_{t}$ gives relatively low weight to motion noise at the frame level and optimizes the model to minimize its impact.
-The rPPG prediction head estimates phase-shifted facial rPPG $\hat{\textbf{y}}_{f}\in R^{t}$ and acral rPPG $\hat{\textbf{y}}_{a}\in R^{t}$, and the input $\boldsymbol{\beta}\in R^{64\times T\times 4\times 4}$ is defined as follows.
+The kernel size and dilation ratio for each convolution layer are denoted by $k$ and $d$ in Fig.
+3.
+The feature extractor consists of the seven atrous convolutions and three max pooling layers and produces a spatiotemporal feature map $\textbf{m}\in\mathbb{R}^{64\times T\times 4\times 4}$.
+In addition, the intermediate feature map $\textbf{m}_{inter}\in\mathbb{R}^{64\times T\times 32\times 32}$ from the feature extractor is utilized in the spatial and temporal attention modules in parallel.
+The spatial attention module consists of three-dimensional convolution, max pooling, average pooling, and pointwise convolution for dimensionality reduction to obtain $\textbf{v}_{s}\in\mathbb{R}^{1\times 1\times 4\times 4}$.
+The spatial attention vector $\boldsymbol{\alpha}_{s}$ is computed by applying the sigmoid function $\sigma$ to $\textbf{v}_{s}$.
+$\boldsymbol{\alpha}_{s}$ gives a higher weight to skin regions to distinguish areas containing physiological information and irrelevant background.
+Similarly, we obtain $\textbf{v}_{t}\in\mathbb{R}^{1\times T\times 1\times 1}$ from the temporal attention module, and acquire the temporal attention vector $\boldsymbol{\alpha}_{t}$.
+$\boldsymbol{\alpha}_{t}$ gives relatively low weight to motion noise at the frame level and optimizes the model to minimize its impact.
+The rPPG prediction heads consist of Siamese structured CNNs and take $\boldsymbol{\beta}$ as an input feature vector, defined as follows.
 $\boldsymbol{\beta}=(\textbf{m}\otimes\boldsymbol{\alpha}_{t})\otimes\boldsymbol{\alpha}_{s},$ (1)
-where $\otimes$ denotes elementwise multiplication, which gives the weights of temporal and spatial attention vectors $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\alpha}_{s}$ to spatiotemporal feature map m.
+where $\otimes$ denotes elementwise multiplication, which broadcasts vectors to a common shape and gives the weights of temporal and spatial attention vectors $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\alpha}_{s}$ to spatiotemporal feature map m.
+Each of the rPPG prediction heads predicts a one-dimensional signal $\hat{\textbf{y}}\in\mathbb{R}^{T}$, representing phase-shifted facial rPPG signals $\hat{\textbf{y}}_{f}$ and acral rPPG signals $\hat{\textbf{y}}_{a}$, respectively.
 Figure 4: Architecture of BBP-Net.
 3.5 Bounded blood pressure network (BBP-Net)
-To estimate SBP and DBP from facial rPPG $\hat{\textbf{y}}_{f}$ and acral rPPG $\hat{\textbf{y}}_{a}$, we designed a multi-scale feature fusion (MSF)-based bounded blood pressure network (BBP-Net), which is shown in Fig.
+We designed a multi-scale feature fusion(MSF)-based bounded blood pressure network (BBP-Net) to estimate SBP and DBP from facial rPPG signal $\hat{\textbf{y}}_{f}$ and acral rPPG signal $\hat{\textbf{y}}_{a}$, as shown in Fig.
 4.
-BBP-Net takes the $6\times T$ set of physiological signals as a stack of velocity plethysmograms (VPGs) and acceleration plethysmograms (APGs), which are the first and second derivatives of the facial and acral rPPG.
-The structure of BBP-Net is designed with a backbone network based on MSF blocks and a parallel BP prediction heads, and scaling sigmoid function.
-The MSF block performs depth-wise separable (DWS) convolution with kernel sizes of the 3 and 5 in parallel.
+BBP-Net takes the physiological signals as a stack of phase-shifted rPPG signals, velocity plethysmography (VPG) and acceleration plethysmography (APG) signals.
+VPG and APG are the first and second derivatives of the facial and acral rPPG signals.
+The structure of BBP-Net is designed with a backbone network based on MSF blocks and parallel BP prediction heads, and scaled sigmoid function.
+$C_{in}$, $C_{out}$, and $C_{mid}$ denote the channels of the input, output, and middle layers of the MSF blocks and the BP prediction heads, respectively, while $T_{in}$ denotes the length of the input signals.
+The MSF block performs depth-wise separable convolution with kernel sizes of the 3 and 5 in parallel and uses Hardswish (Howard et al., 2019) as an activation function.
 Then it performs a linear operation and a channel-wise softmax operation to compute the weight vector.
-Every MSF blocks contain the residual connections.
+All of MSF blocks contain the residual connections.
 The BP prediction head includes the computation process to obtain the weight vector by employing the bottleneck attention module (BAM) (Park et al., 2018) and residual connections in parallel.
-We modify the sigmoid function into a scaling sigmoid function to bound the estimated blood pressure to the normal range, and it is defined as follows.
-$\hat{y}_{BP}=BP_{min}+\frac{BP_{max}-BP_{min}}{1+exp({-z}+\tau)},\,where\,BP\in\left\{SBP,DBP\right\},$ (4)
-where $\hat{y}_{BP}$ is the estimated blood pressure, and ${BP}_{max}$ and ${BP}_{min}$ are heuristic parameters that represent the range of reasonable blood pressure.
-$z$ and $\tau$ are the logit estimated from the BP prediction head and temperature, respectively.
+We modify the sigmoid function into a scaled sigmoid function to bound the estimated blood pressure to the predefined range, and it is defined as follows.
+$\widehat{BP}=BP_{min}+(BP_{max}-BP_{min})\sigma(z/\tau),\,$ (2)
+where $\widehat{BP}$ is the estimated blood pressure, and ${BP}_{max}$ and ${BP}_{min}$ are heuristic parameters that represent the range of reasonable blood pressure.
+The symbol $BP$ consists of two types of blood pressure, $SBP$ and $DBP$, and the predefined bounded ranges of $\widehat{SBP}$ and $\widehat{DBP}$ are set to [155, 85] and [95, 45] mmHg, respectively.
+$\sigma$, $z$ and $\tau$ denote the sigmoid function, the logit estimated from the BP prediction head and temperature, respectively.
+We set temperature to 2 in our experiments.
 The structure of BBP-Net is based on the architecture proposed by  Hu et al.
-(2022), which is designed to estimate SBP and DBP from PPG, and has achieved state-of-the-art (SOTA) performance on the UCI-dataset.
-While previous studies that take a single PPG of size $3\times T$ as input, we utilize the phase-shifted information of the rPPG and employ a scaling sigmoid function.
+(2022), which is designed to estimate SBP and DBP from PPG signals.
+While previous studies that take a stack of PPG, VPG, APG signals as input, we utilize the phase-shifted information of the rPPG signals and employ a scaled sigmoid function.
 3.6 Loss functions
-Since the purpose of DRP-Net and BBP-Net is regression to estimate the biometrics of rPPG, heart rate, and blood pressure, we designed a loss function based on the mean squared error (MSE).
-The loss function $L_{HR}$ of DRP-Net for heart rate estimation is denoted as
-$L_{HR}=\frac{1}{n}\sum_{k=1}^{n}\left(\hat{y}_{HR,k}-y_{HR,k}\right)^{2},$ (5)
-where $\hat{y}_{HR,k}$ and $y_{HR,k}$ are the predicted and ground truth heart rate, respectively, and $n$ is the number of data samples.
-The loss function in the frequency domain $L_{frequency}$ is defined based on the power spectrum of the physiological signal via FFT and can be defined as follows.
-$L_{frequency}=\frac{1}{n}\sum_{k=1}^{n}\left(FFT(\hat{\textbf{y}}_{k})-FFT(\textbf{y}_{PPG,k})\right)^{2},$ (6)
-where $\hat{\textbf{y}}_{k}$ and $\textbf{y}_{PPG,k}$ denote the rPPG predicted by DRP-Net and the PPG obtained from the sensor, respectively.
-In the case of using only frequency domain losses, the range of predicted signal’s amplitude is relatively small, therefore we propose $L_{peakvalley}$ as a loss function for scaling.
-$L_{peakvalley}=\frac{1}{n}\sum_{k=1}^{n}\left[\left\{\phi(\hat{\textbf{y}}_{k}^{P})-\phi(\textbf{y}_{PPG,k}^{P})\right\}^{2}+\left\{\phi(\hat{\textbf{y}}_{k}^{V})-\phi(\textbf{y}_{PPG,k}^{V})\right\}^{2}\right],$ (7)
-where $P$ and $V$ denote the peak and valley values of the signal, respectively, and the average of these values is denoted by $\phi$.
-$L_{peakvalley}$ is the time domain loss function but it is independent of the phase of the predicted signal.
-In addition, we define a loss function $L_{temporal}$ in the time domain to train an acral rPPG that simulates the scale and phase of the PPG.
-It is defined as
-$L_{temporal}=\frac{1}{n}\sum_{k=1}^{n}\left(\hat{\textbf{y}}_{k}-\textbf{y}_{PPG,k}\right)^{2}.$ (8)
-The total loss function to optimize the facial rPPG is $L_{facial}=\lambda_{1}L_{HR}+\lambda_{2}L_{frequency}+L_{peakvalley}$, which guides the model to distinguish pulse waves in the facial image sequence.
-On the other hand, acral rPPG utilizes $L_{acral}=\lambda_{1}L_{HR}+\lambda_{2}L_{frequency}+L_{peakvalley}+L_{temporal}$ to simulate the phase of the PPG.
-BBP-Net for blood pressure estimation employs Huber loss (Collins, 1976) to predict SBP and DBP which we define as $L_{BP}$.
-Additionally, acral rPPG scaled with SBP and DBP ranges is adjusted by minimizing $L_{ABP}$ to the ABP signal.
+The loss functions for training DRP-Net are $L_{freq}$ and $L_{HR}$ in the frequency domain, and $L_{time}$ and $L_{pv}$ in the time domain.
+The frequency domain loss $L_{freq}$ is designed based on the PSD of the physiological signals and defined as
+$L_{freq}=\lVert P(\hat{\textbf{y}})-P(\textbf{y})\lVert_{2},$ (3)
+where $\hat{\textbf{y}}$ and y denote the rPPG signals predicted by DRP-Net and the pseudo PPG signals obtained from the ABP signals, respectively.
+$P(\cdot)$ indicates the PSD of the signals, which produces a real-valued vector obtained after applying the fast Fourier transform.
+The frequency range of the vectors is between 0.5 Hz and 3 Hz, corresponding to heart rates between 30 and 180 BPM.
+Moreover, the heart rate corresponds to the frequency with the maximum amplitude in the PSD, indicating the distinct period in the physiological signal.
+The loss function of heart rate $L_{HR}$ is defined as follows,
+$L_{HR}=\lVert\widehat{HR}-HR\rVert_{2},$ (4)
+where $\widehat{HR}$ and $HR$ are the predicted and ground truth heart rate, respectively.
+In addition, we define the time domain loss $L_{time}$ to estimate the acral rPPG signal, which mimics the scale and phase of the pseudo PPG signal, with the following equation (5).
+$L_{time}=\lVert\hat{\textbf{y}}-\textbf{y}\rVert_{2}.$ (5)
+We propose an additional time domain loss $L_{pv}$, which is phase-independent and designed to align the scale of facial rPPG signals.
+To remove noise and improve smoothness, we performed standardization, detrending, and filtering on the rPPG signals.
+We extract all of temporal location where the first derivative of the signal changes from positive to negative to detect all peaks in the window.
+This process is defined as follows,
+$\tilde{\textbf{D}}^{r}(s_{t})=\left\{s_{t}\mid(s_{t}-s_{t-1})(s_{t+1}-s_{t})<0\right\}\left\{\begin{matrix}r=p~{}~{}if~{}s_{t}>s_{t-1},\\ r=v~{}~{}if~{}s_{t}<s_{t-1,}\end{matrix}\right.$ (6)
+where $\tilde{\textbf{D}}^{r}(\cdot)$, $s$, $t$, $p$, and $v$ denote set of detected peaks or valleys, input signal, sample of signal, peak and valley, respectively.
+The values lower than the average of detected peaks are removed to minimize the effect of the dicrotic notch in each pulse.
+The remaining peaks indicate a vector composed of the maximum value of each pulse.
+The loss is calculated as the average of the remaining peaks using the following equation (7), considering temporal discrepancy and difference of vector sizes between rPPG and pseudo PPG signals.
+$D^{p}(s)=\left\{\Phi(\tilde{\textbf{D}}_{i}^{p}(s))\mid\tilde{\textbf{D}}_{i}^{p}(s)>\Phi(\tilde{\textbf{D}}^{p}(s))\right\},$ (7)
+where $D^{p}(\cdot)$, $\Phi(\cdot)$, and $i$ denote peak value within the window, average, and $i$th sample of detected peaks, respectively.
+In the case of valleys, the standards for detection and removal are opposite to those for peaks, as defined as follows,
+$D^{v}(s)=\left\{\Phi(\tilde{\textbf{D}}_{i}^{v}(s))\mid\tilde{\textbf{D}}_{i}^{v}(s)<\Phi(\tilde{\textbf{D}}^{v}(s))\right\}.$ (8)
+$L_{pv}$ is defined based on the L2 norm of the peak valley matrix of the predicted rPPG and pseudo PPG signals, with the following equation,
+$L_{pv}=\lVert\hat{\textbf{c}}-\textbf{c}\rVert_{2},~{}where~{}\hat{\textbf{c}}=[D^{p}(\hat{y}),D^{v}(\hat{y})],\textbf{c}=[D^{p}(y),D^{v}(y)],$ (9)
+where $\hat{\textbf{c}}$ and c denote the peak valley matrix of predicted rPPG and pseudo PPG signals, respectively.
+The total loss function to optimize the facial rPPG signals is defined as
+$L_{facial}=\lambda_{1}L_{HR}+\lambda_{2}L_{freq}+L_{pv},$ (10)
+which guides the model to extract pulse waves in the facial image sequence. On the other hand, optimization of acral rPPG signals utilizes the following equation,
+$L_{acral}=\lambda_{1}L_{HR}+\lambda_{2}L_{freq}+L_{pv}+L_{time},$ (11)
+to mimic the phase of the pseudo PPG signal which obtained from acral site. We experimented by setting the constant $\lambda_{1}$ and $\lambda_{2}$ to 0.0001 and 100, respectively.
+We designed $L_{BP}$ and $L_{ABP}$ as loss functions of BBP-Net for blood pressure estimation.
+$L_{BP}$ employs Huber loss (Collins, 1976) to calculate the loss for the predicted SBP and DBP, defined by the following equation,
+$L_{BP}=\left\{\begin{matrix}\frac{1}{2}(\widehat{BP}-BP)^{2}&if\left|(\widehat{BP}-BP)\right|<\delta,\\ \delta\left|(\widehat{BP}-BP)\right|-\frac{1}{2}\delta),&otherwise,\end{matrix}\right.$ (12)
+where $\widehat{BP}$ and $BP$ denote for predicted and ground truth blood pressure and heuristic parameter $\delta$ which set to 1 in the experiment, determines the type of loss function for the error.
+The Huber loss is based on the L2 and L1 norm, applying the former for errors within a defined $\delta$ range for smoothness and differentiability, and the latter for larger errors to enhance robustness against outliers.
+In addition, we reconstructed the ABP signal in the time domain using the predicted physiological information, scaling $\hat{\textbf{y}}_{a}$ to the range of $\widehat{SBP}$ and $\widehat{DBP}$ as follows,
+$\hat{\textbf{y}}_{s}=\frac{\hat{\textbf{y}}_{a}-\hat{y}_{min}}{\hat{y}_{max}-\hat{y}_{min}}(\widehat{SBP}-\widehat{DBP})+\widehat{DBP},$ (13)
+where $\hat{y}_{max}$ and $\hat{y}_{min}$ are the maximum and minimum values of $\hat{\textbf{y}}_{a}$, and $\hat{\textbf{y}}_{s}$ denotes scaled acral rPPG signal.
+$L_{ABP}$ is defined based on the L2 norm of scaled acral rPPG signal $\hat{\textbf{y}}_{s}$ and the ABP signal Y, defined as follows,
+$L_{ABP}=\lVert\hat{\textbf{y}}_{s}-\textbf{Y}\rVert_{2}.$ (14)
+$L_{ABP}$ fine-tunes $\widehat{SBP}$ and $\widehat{DBP}$ to exist within the ranges of peak and valley of the ABP signal as they change from pulse to pulse.
 4 Experimental results
@@
 Python and Pytorch were used for the implementation of the proposed algorithm.
-For experimental details, a time window of 150 (6 seconds), a batch size of 8, and a decay factor of 0.99 were used, and the learning rate of the Adam optimizer was set to 0.001 and 0.0001 for MMSE-HR and V4V, respectively.
-To evaluate the accuracy of estimated heart rate and blood pressure, we adopted the mean absolute error (MAE) and root mean squared error (RMSE).
-Moreover, the proposed method was evaluated using the Pearson correlation coefficient r.
+For experimental details, a window of 150 (6 seconds), a batch size of 8, and a decay factor of 0.99 were used, and the learning rate of the Adam optimizer was set to 0.001 and 0.0001 for MMSE-HR and V4V, respectively.
+To evaluate the accuracy of estimated heart rate and blood pressure, we adopted the mean absolute error (MAE), root mean squared error (RMSE), Pearson correlation coefficient r.
 4.1 Datasets
-To demonstrate the heart rate and blood pressure estimation accuracy of the proposed method, we experimented on the MMSE-HR (Multimodal Spontaneous Expression-Heart Rate) (Zhang et al., 2016b) and V4V (Vision for Vitals) database (Revanur et al., 2021).
-The two datasets are sub-datasets of the MMSE database (BP4D+), which consists of synchronized facial image sequences and continuous blood pressure signals.
-As introduced in Section 3.1, the biometric information of PPG, heart rate, SBP, and DBP can be calculated by preprocessing the blood pressure signals.
-To ensure data diversity, tasks were performed to arouse amusement, pain, anger, and disgust.
-RGB video was recorded at the resolution of 1040$\times$1392 and the frame rate of 25 FPS, and blood pressure signals were collected at the sampling rate of 1000 Hz.
+We conducted our experiment using the MMSE-HR (Multimodal Spontaneous Expression-Heart Rate) (Zhang et al., 2016b) and V4V (Vision for Vitals) database (Revanur et al., 2021) to demonstrate the heart rate and blood pressure estimation performance of the proposed method.
+The utilized datasets are sub-datasets derived from the MMSE database (BP4D+), which consists of synchronized facial image sequences and continuous ABP signals.
+As introduced in Section 3.1, the physiological information of pseudo PPG signals, heart rate, SBP, and DBP can be calculated from ABP signals.
+For physiological diversity in the dataset, subjects were conducted various tasks to arouse amusement, pain, anger, and disgust.
+RGB video was recorded at the resolution of 1040$\times$1392 with a frame rate of 25 FPS, and ABP signals were collected at the sampling rate of 1000 Hz.
 The MMSE-HR database released at the CVPR 2016 for The Heart Rate Estimation Algorithm Challenge includes 102 samples from 17 male and 23 female subjects.
-The average signal lengths are 30 seconds.
-The V4V database released at the ICCV workshop 2021 for The 1st Vision for Vitals Challenge includes 1,358 data samples split into 724 training, 276 validation, and 358 test samples.
-The average signal lengths are 40 seconds, collected from 179 subjects.
+The average length of signals is 30 seconds.
+The V4V database released at the ICCV workshop 2021 for The 1st Vision for Vitals Challenge includes 1,358 data samples from 179 subjects split into 724 for training, 276 for validation, and 358 for test samples.
+The average length of signals is 40 seconds.
 Table 1: Comparison of heart rate estimation results on the MMSE-HR database.
@@
 Ours (acral) 6 1.91 4.74 0.93
-Figure 5: Pearson correlation graph of predicted and actual HR
-Figure 6: Predicted facial(blue), acral(green) rPPG and PPG(red)
+Figure 5: Pearson correlation graph of predicted and actual heart rate
+Figure 6: Predicted facial (blue), acral (green) rPPG and pseudo PPG (red) signals
 4.2 Quantitative results on MMSE-HR database
-In this section, we compare the heart rate estimation performance of the proposed algorithm with previous methods using the public dataset MMSE-HR.
-To deal with the size of the MMSE-HR database, we employed the 5-fold cross-validation method for heart rate estimation, and the folds were split independently between subjects to demonstrate generalizability.
+In this section, we compare the heart rate estimation performance of the DRP-Net with previous methods using the public dataset MMSE-HR.
+To deal with the small size of the MMSE-HR database, we employed the 5-fold cross-validation method for heart rate estimation, and the folds were split independently between subjects to demonstrate generalizability.
 Table.
-1 presents the average heart rate estimation performance of facial and acral rPPG.
-The proposed method achieved MAE, RMSE, and r of 1.85, 4.51, and 0.948, respectively, and outperformed previous deep learning-based approaches.
-The Pearson correlation of predicted and actual heart rate based on facial rPPG per fold is shown in Fig.
+1 presents the average of fold-wise heart rate estimation performance of facial and acral rPPG signals.
+The proposed method achieved MAE, RMSE, and r of 1.78, 4.27, and 0.95, respectively, and outperformed previous deep learning-based approaches.
+The Pearson correlation of actual and predicted heart rate based on facial rPPG signals is shown in Fig.
 5.
-This graphs represent a linear trend, indicating a significant similarity between the predicted and actual heart rate.
+These graphs show a linear trend, indicating a significant similarity between the predicted and actual heart rate.
 Algorithms (Nowara et al.
 (2021), Jaiswal & Meenpal (2022), Ouzar et al.
-(2023)) with short input image sequences provide fast inference times but at the cost of higher error rates.
+(2023)) with short input image sequences provide fast inference times.
+However, the loss of temporal information leads to an increase in estimation error.
 Video transformer-based algorithms such as Yu et al.
 (2022) and Yu et al.
-(2023) show considerable performance in cross-dataset tests.
-However, they require high-performance computing equipment due to their extensive parameters and computational demands.
+(2023) demonstrate significant performance in cross-dataset tests.
+However, transformers require high-performance computing resources due to their extensive parameters and computational demands.
 Additionally, Fig.
-6 shows the 2 predicted rPPG and their phase differences as the cardiac cycle increases.
-Comparing the 2 approaches in our proposed method in Table.
-1, it indicates that the heart rate estimation error based on the facial rPPG is lower than the acral rPPG.
-These experimental findings suggest that forcing the rPPG phase leads to decreased model performance and emphasize the importance of inferring actual pulse waves from captured region.
+6 shows the predicted facial and acral rPPG signals and their temporal discrepancy.
+Comparing the two approaches in our proposed method in Table.
+1, the results indicate that the heart rate estimation error based on the facial rPPG signals is lower than the acral rPPG signals.
+These experimental findings suggest that forcing the phase of rPPG signal leads to decreased model performance and emphasize the importance of inferring actual pulse waves from a captured region.
 Table 2: Comparison of blood pressure estimation results on the MMSE-HR database.
@@
 Table.
-2 presents the results of the blood pressure estimation experiment on the MMSE-HR database, and Fig.
-7 shows the results of the Bland-Altman plot.
-The Rong & Li (2021) and Schrumpf et al.
-(2021) algorithms extracted rPPG using a mathematical approach and estimated blood pressure through handcrafted features and deep learning-based models, respectively.
-These studies are similar to our approach in extracting rPPG as an intermediate step.
-However, our method improves performance by considering phase-shifted information for blood pressure estimation.
+2 presents the results of the blood pressure estimation experiment on the MMSE-HR database.
+Fig.
+7 shows the blood pressure estimation results of the Bland-Altman plots, and the solid line and dotted lines represents mean of difference and 95% limits of agreement, respectively.
+The algorithms proposed by Rong & Li (2021) and Schrumpf et al.
+(2021) extracted rPPG signals using a non-parametric approach and estimated blood pressure via handcrafted feature-based and deep learning-based models, respectively.
+These studies are similar to our approach in extracting rPPG signals as an intermediate step.
+However, the difference of our method lies in the utilization of phase-shifted rPPG signals derived from a parametric model.
 Chen et al.
-(2023) proposed a blood pressure estimation model based on a 2 dimensional spatiotemporal map, a video preprocessing technique based on the YUV color space.
-However, they neglect pulse waves at the acral sites of the body, resulting in higher SBP and DBP error rates compared to the proposed method.
+(2023) proposed a blood pressure estimation model that utilizes a two-dimensional spatiotemporal map, a video preprocessing technique based on various color spaces.
+However, they neglect to consider temporal discrepancy from the cardiac cycle, leading to increased errors in estimating SBP and DBP.
 The experimental results in Table.
-2 represent that the temporal information in the pulse wave is meaningful for non-contact blood pressure estimation.
+2 represent that the temporal discrepancy in the rPPG signals is meaningful for non-contact blood pressure estimation.
 4.3 Quantitative results on V4V database
@@
 Ours (acral) 4.13 10.14 0.725
-In this section, we conducted experiments on V4V database to compare the heart rate estimation performance of the proposed DRP-Net with existing algorithms and the experimental results are shown in Table.
-3.
-The experiments are conducted on the training and test data, and the graph 6 in Fig.
-5 shows a Pearson correlation plot of the predicted and actual heart rates.
-The metrics for DRP-Net heart rate estimation performance using V4V are the same as for the MMSE-HR database.
-The proposed method using face rPPG outperformed in MAE and r, with 3.83 and 0.75, respectively.
-Additionally, the proposed method with acral rPPG achieved MAE, RMSE, and r of 4.13, 10.14, and 0.72, respectively This result is a slightly higher error than facial rPPG, similar to the results on the MMSE-HR database.
-The performance comparison between Yu et al.
-(2019) and DRP-Net shows that the wider receptive fields through 3D atrous convolution is beneficial for heart rate estimation than the encoder-decoder structure based on 3D CNN.
-The performance comparison with Kim et al.
-(2022) shows contrasting MAE and RMSE results.
-It indicates that the outliers of heart rate predicted by the proposed method are larger, while the average error is smaller.
 Table 4: Comparison of blood pressure estimation results on the V4V database.
@@
 Ours 13.640 16.782 9.402 11.903
+In this section, we present experiments conducted on the V4V database to compare the heart rate estimation performance of the proposed DRP-Net with existing algorithms, as shown in Table 3.
+Facial rPPG signals derived from the proposed method outperformed others in terms of MAE and r, with values of 3.83 and 0.75, respectively.
+The sixth graph in Fig.
+5 shows a Pearson correlation plot of the predicted and actual heart rates.
+Additionally, the proposed method with acral rPPG signals achieved MAE, RMSE, and r of 4.13, 10.14, and 0.72, respectively.
+This result indicates a slightly higher error compared to the results of facial rPPG signals, consistent with findings from the MMSE-HR database.
+The performance comparison between Yu et al.
+(2019) and DRP-Net indicates that utilizing wider receptive fields through atrous convolution is more advantageous for learning spatiotemporal features than the encoder-decoder structure based on three dimensional CNN.
+The performance comparison with Kim et al.
+(2022) shows contrasting MAE and RMSE results.
+This result indicates that the error of outliers of proposed method is large.
 The blood pressure estimation results of BBP-Net is shown in Table.
 4.
-Our proposed algorithm achieved MAE and RMSE of 13.64, 16.78 and 9.40, 11.90 for SBP and DBP, respectively.
+Our proposed algorithm achieved MAE, RMSE of 13.64, 16.78 and 9.40, 11.90 for SBP and DBP, respectively.
+Blood pressure estimation performance of BBP-Net outperforms baseline algorithms such as AlexNet, ResNet, and LSTM, as well as Schrumpf et al.
+(2021) and the algorithms proposed by Hamoud et al.
+(2023).
 The performance of the baseline algorithms (AlexNet, ResNet, LSTM) is reported in Hamoud et al.
 (2023).
-Blood pressure estimation performance of BBP-Net outperforms baseline algorithms such as AlexNet, ResNet, and LSTM, as well as Schrumpf et al.
-(2021) and the algorithm proposed by Hamoud et al.
-(2023).
 While the SBP and DBP estimation performance presented in the Hamoud et al.
-(2023) has the limitation that each performance is a combination of different models, the BBP-Net estimates both blood pressures simultaneously.
-As shown in the experimental results on the V4V database, the rPPG estimated from the 2 body parts and its phase-shifted information are effective in improving the estimation performance of blood pressure.
+(2023) has the limitation that each performance is a combination of different models, the BBP-Net estimates both of blood pressures simultaneously.
+As shown in the experimental results on the V4V database, the rPPG signals estimated from the two physiological sites and its phase-shifted information are effective in improving the estimation performance of blood pressure.
 4.4 Ablation study
-Table 5: Comparative results of mean absolute error in heart rate estimation task across various window sizes and folds.
+Table 5: Comparative results of heart rate estimation performance across various window lengths and folds.
 Window
@@
 acral 1.13 1.04 0.58 1.45 3.20 1.48
-Table 6: Comparison of MAE in heart rate estimation with and without data augmentation.
-Augmentation rPPG MAE (BPM)
-X facial 2.149
+Table 6: Comparative results of heart rate estimation performance with data augmentation and $L_{pv}$.
+Augmentation $L_{pv}$ rPPG MAE (BPM)
+facial 2.149
 acral 2.002
-O facial 1.783
+facial 3.670
+acral 4.178
+facial 1.783
 acral 1.909
-Table 7: Comparison of MAE in heart rate estimation with and without peak valley loss.
-Peak valley loss rPPG MAE (BPM)
-X facial 3.670
-acral 4.178
-O facial 1.783
-acral 1.909
-The window size of the input image sequence is an important heuristic parameter in terms of the computational power and the latency of model.
-Smaller window sizes have the advantage of lower latency in inference time, while the loss of temporal information can decrease the performance.
-Table.
-5 presents the performance with different window size and fold, and it demonstrates that the average heart rate estimation performance increases as the window size increases.
-The reduced performance observed at a window size of 2 seconds is due to the dilation ratio of DRP-Net being relatively large compared to the sequence information.
-However, estimation of blood pressure requires long temporal information of rPPG, therefore we selected window size of 6 seconds and a dilation ratio of 4.
-Table.
-6 indicates the results of heart rate estimation of DRP-Net with and without data augmentation techniques.
-The performance of all estimated rPPG improved after data augmentation.
-This result demonstrate that training with data from bradycardia and tachycardia via frame interpolation and down-sampling improves the generalization performance of the model.
-In addition, the experimental results of excluding the peak and valley guides from the loss function are demonstrated in Table.
-7 to investigate the impact of $L_{peakvalley}$.
-The average MAE without $L_{peakvalley}$ increases about 2 BPM, which indicates that the amplitude information in the PPG contributes the improvement of heart rate estimation performance.
-Table 8: Comparative results of MAE in blood pressure estimation for the input rPPG types.
+Table 7: Comparative results results of blood pressure estimation performance for the type of rPPG.
 Input rPPG SBP MAE (mmHg) DBP MAE (mmHg)
@@
 both 10.192 7.090
-Table 9: Comparison of blood pressure estimation performance with and without the scaling sigmoid function.
-Scaling sigmoid function SBP DBP
+Table 8: Comparative results of blood pressure estimation performance with and without the scaled sigmoid function.
+Scaled sigmoid function SBP DBP
 MAE (mmHg) RMSE MAE (mmHg) RMSE
@@
 O 10.19 13.01 7.09 8.86
-The ablation study for the blood pressure estimation model consists of the types of input rPPG and the existence of the scaling sigmoid function.
+The ablation study on heart rate estimation is conducted to investigate the impact of the proposed methods.
+The window length of the input image sequence is an important heuristic parameter in terms of the computational power and the latency of the model.
+Smaller window length have the advantage of lower latency in inference time, however, the loss of temporal information can decrease performance.
 Table.
-8 presents the estimation error of the SBP and DBP based on the facial or acral rPPG, and the estimation performance of both cases are lower than using dual rPPG.
-This results indicate that the phase-shifted information of pulse waves in different body parts contributes to improved estimation performance of SBP and DBP.
-Additionally, Table.
-9 shows the performance of blood pressure estimation with and without the scaling sigmoid function.
-Without the scaling sigmoid function, the MAE for SBP and DBP was significantly increased to 9 mmHg and 3 mmHg, respectively.
+5 presents the performance with different window length, and demonstrates that the heart rate estimation performance increases as the window length increases.
+The lowest performance observed at a window length of 2 seconds is due to the dilation ratio of DRP-Net being relatively large compared to the sequence length.
+However, estimation of blood pressure requires long-term information of physiological signals, therefore we selected a window length of 6 seconds and a dilation ratio of 4.
+Table.
+6 indicates the results of heart rate estimation of DRP-Net with and without data augmentation techniques and $L_{pv}$.
+The performance of rPPG estimation improved after data augmentation.
+This result demonstrate that training with data from bradycardia and tachycardia via frame interpolation and down-sampling improves the generalizability.
+Additionally, the experimental results of excluding the peak and valley guides from the loss function are also demonstrated in Table.
+6 to investigate the impact of $L_{pv}$.
+The average MAE with $L_{pv}$ decreases by about 2 BPM, which indicates that the guidance of scale significantly affects the estimation of accurate heart rate.
+The ablation study for BBP-Net focuses on the input types of rPPG and the utilization of the scaled sigmoid function.
+Table 7 presents the estimation errors for blood pressure from different input types of rPPG signals and shows the superior performance of phase-shifted rPPG.
+This results indicate that the temporal discrepancy of pulse waves in different physiological sites contributes to improved estimation performance of blood pressure.
+Table.
+8 shows the performance of blood pressure estimation with and without the scaled sigmoid function of BBP-Net.
+Without the scaled sigmoid function, the MAE for SBP and DBP increased by 9 mmHg and 3 mmHg, respectively.
 These experimental results suggest the importance of guiding the model towards a reasonable blood pressure range.
-4.5 International standards and blood pressure estimation performance
-Table 10: Comparison results of blood pressure estimation performance with BHS standards.
+4.5 International standards
+Table 9: Comparison of blood pressure estimation performance with BHS standards.
 $\leq$ 5 mmHg $\leq$ 10 mmHg $\leq$ 15 mmHg
@@
 grade C 40 % 65 % 85 %
-Table 11: Comparison results of blood pressure estimation performance with AAMI standards.
+Table 10: Comparison of blood pressure estimation performance with AAMI standards.
 ME (mmHg) SDE (mmHg)
@@
 AAMI standard $\leq$5 $\leq$8
-We compare the performance of our proposed blood pressure estimation algorithm with international standards, the British Hypertension Society (BHS)  (O’Brien et al., 1990) and The Association for the Advancement of Medical Instrumentation (AAMI)  (Stergiou et al., 2018).
-The BHS standard evaluates the absolute number of blood pressure estimation errors as a cumulative percentage, and Table.
-10 presents the standards and blood pressure estimation performance.
+We compare the performance of proposed algorithm with international standards, the British Hypertension Society (BHS)  (O’Brien et al., 1990) and The Association for the Advancement of Medical Instrumentation (AAMI)  (Stergiou et al., 2018).
+The BHS standard evaluates the absolute number of blood pressure estimation errors as a cumulative percentage, as shown in Table.
+9.
 On the MMSE-HR database, the proposed method achieves grade B with 81 % DBP estimation performance for errors of 5 mmHg or less.
-Additionally, DBP estimation performance on the V4V database also reached grade C.
+Additionally, DBP estimation performance on the V4V database also achieved grade C.
 The AAMI standard in Table.
 10 indicates that blood pressure estimation performance meets the standard when the mean error (ME) and standard deviation error (SDE) are under 5 and 8 mmHg, respectively.
-In the MMSE-HR database, the estimation performance of SBP and DBP for ME is -0.13 mmHg and 1.58 mmHg, respectively, which meets the standard.
-The result on the V4V database shows that ME is 5.96 mmHg at SBP and -2.4 mmHg at DBP.
+In the MMSE-HR database, the blood pressure estimation performances in terms of ME are -0.13 mmHg and 1.58 mmHg, respectively, which meet the standard.
 5 Conclusion
-This paper proposed the deep learning framework based on facial images for contactless heart rate and blood pressure estimation.
-We assumed that phase differences exist between pulse waves due to inconsistencies in data measurement sites.
-Based on this assumption, we improved blood pressure estimation performance and demonstrated with extensive experiments.
-To estimate pulse waves of 2 physiological sites from facial images, we proposed the DRP-Net that focuses on spatiotemporal feature extraction.
-DRP-Net estimates facial rPPG using information in the time-global frequency domain to predict pulse waves from the face, and predicts acral rPPG that mimic the phase of PPG.
-As a result, the proposed model achieved the MAE of 1.78 and 1.91 BPM based on acral and facial rPPG in heart rate estimation task.
-Moreover, we proposed BBP-Net to estimate blood pressure based on the phase analysis and temporal features of the dual rPPG, applying the modified sigmoid function to bound the predicted blood pressure to the reasonable range.
-In the blood pressure estimation task, we achieved MAE of SBP and DBP of 10.19 mmHg and 7.09 mmHg, respectively.
-However, the public datasets used in this study have the limitation in that they do not include pulse wave data from the face, making it unable to verify the accuracy of the predicted facial rPPG phase.
-We were also constrained by the absence of other public datasets containing both facial images and blood pressure data, limiting experimentation in various domains.
-Inspired by these limitations, we have plan to collect real-world dataset which includes facial videos, PPG, and blood pressure with various data collection scenarios and develop sensor acquisition system in our future research.
-Additionally, we will conduct research on improving the performance of physiological measurement system using the collected real-world dataset.
+This paper proposes a deep learning framework based on facial images for contactless heart rate and blood pressure estimation.
+We assumed temporal discrepancy between pulse waves due to potential inconsistencies in data measurement sites.
+To estimate phase-shifted rPPG signals, we propose DRP-Net which learns spatiotemporal features from facial images via frequency and time domain loss functions.
+Additionally, frame sampling based augmentation is proposed to augment bradycardia and tachycardia data.
+Moreover, we propose BBP-Net to estimate blood pressure from phase-shifted rPPG signals and its temporal discrepancy.
+In the experimental results, we demonstrated the improvement in heart rate and blood pressure estimation performance by previous methods and ablation studies.
+However, we found limitations in most public datasets, which exclude facial PPG signals or blood pressure data.
+Inspired by founded limitations, we are planning to collect a real-world dataset with various scenarios to verify phase of facial rPPG signals and the performance of the proposed deep learning framework.
+Finally, we hope that our experimental findings can be actively utilized in further research for robust remote physiological measurement.
+Acknowledgements
+Collate acknowledgements in a separate section at the end of the article before the references and do not, therefore, include them on the title page, as a footnote to the title or otherwise.
+List here those individuals who provided help during the research (e.g., providing language help, writing assistance or proof reading the article, etc.).